### v1

# Should we pre-train a decoder in contrastive learning for dense prediction tasks?

Abstract
Contrastive learning in self-supervised settings primarily focuses on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. This conventional approach, however, overlooks the potential benefits of jointly pre-training both the encoder and decoder. In this paper, we propose DeCon: a framework-agnostic adaptation to convert an encoder-only self-supervised learning (SSL) contrastive approach to an efficient encoder-decoder framework that can be pre-trained in a contrastive manner. We first update the existing architecture to accommodate a decoder and its respective contrastive loss. We then introduce a weighted encoder-decoder contrastive loss with non-competing objectives that facilitates the joint encoder-decoder architecture pre-training. We adapt two established contrastive SSL frameworks tailored for dense prediction tasks, achieve new state-of-the-art results in COCO object detection and instance segmentation, and match state-of-the-art performance on Pascal VOC semantic segmentation. We show that our approach allows for pre-training a decoder and enhances the representation power of the encoder and its performance in dense prediction tasks. This benefit holds across heterogeneous decoder architectures between pre-training and fine-tuning and persists in out-of-domain, limited-data scenarios.
**footnotetext: Equal contributions for this work. Co-first authors are ordered in alphabetical order.
1 Introduction
The increasing demand for efficient Deep Learning (DL) methods arises from their ability to address complex tasks. While traditional approaches like supervised learning can yield powerful, high-performing models, they depend heavily on large volumes of high-quality labeled data. However, obtaining such annotations is often tedious, expensive, and sometimes impractical, even when abundant data is available. To overcome this bottleneck, one can benefit from a pre-trained model. Pre-trained models are usually models trained from scratch on a large scale dataset and allow for better performance on downstream tasks with smaller datasets. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on part of or on a full annotated dataset. ImageNet [15] is a large and widely utilized dataset with annotated classes. However, models pre-trained on ImageNet classification are not the best fit for every downstream tasks, especially when the downstream objective is not classification. Since creating large-scale annotated dataset for every tasks presents numerous challenges, Self-Supervised Learning (SSL) has emerged as a promising alternative, enabling models to learn from the vast amount of unlabeled data and perform effectively on various downstream tasks. These pre-trained models can reach downstream performance close to a fully supervised approach using much smaller to no labeled datasets. Notable approaches such as SimCLR [7], VicReg [2], and MoCo [22] showed promising results by leveraging contrastive learning for downstream classification. However, most SSL frameworks primarily target classification downstream tasks. Some efforts such as PixCon [37] or DenseCL [45], focus on dense prediction tasks like segmentation or object detection. In all of these efforts, authors only train the encoders of the DL architectures and focus on incorporating a local loss with existing classification-targeted SSL approaches to adapt with dense tasks.
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, a framework agnostic encoder-decoder adaptation to pre-train an encoder jointly with a decoder, improving its representation power and efficiently preparing it to be combined with a decoder in downstream tasks. Our contributions are as follows.
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder with a decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even without transferring the pre-trained decoder.
  • We improve on this adaptation by introducing DeCon-ML, which extends DeCon-SL with a multi-level decoder loss function and an encoder level-wise channel dropout to promote a comprehensive feature usage of the encoder. It achieves state-of-the-art results on COCO object detection and instance segmentation and matches top performance on Pascal VOC semantic segmentation.
  • We demonstrate that DeCon improves the encoder’s downstream performance compared to encoder-only frameworks, even when pre-training and fine-tuning decoders differ.
  • We show that DeCon’s efficiency persists in out-of-domain downstream tasks and provides even greater benefits in scenarios with limited annotated fine-tuning data.
2 Related works
Self-Supervised pre-training: In SSL, deep learning models are trained on pretext tasks where the data itself provides supervision, which allow models to learn meaningful representations without any manual annotations. Early SSL works focused on designing these tasks, such as predicting missing parts of an image (image inpainting) [38], context prediction [16], solving jigsaw puzzles [34], colorizing grayscale images [52], and predicting rotations [20]. These methods aimed to create surrogate objectives that encourage models to develop feature representations, which could later be transferred to downstream tasks. However, these methods often produced task-specific representations with limited generalizability and restricted scalability.
Recent methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches in SSL aim to learn meaningful representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 51], or adversarial methods, where a generative model learns to synthesize realistic samples while an encoder captures high-level features [48, 26]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—negative pairs [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum contrast [22] or other architectural and optimization techniques are employed [5].
Dense pre-training: Even though most of the aforementioned frameworks can be used for downstream segmentation or detection, these models are not designed specifically for pixel-level, i.e., dense downstream tasks. Consequently, developing SSL approaches that generalize well to dense downstream tasks, such as the segmentation task, is an active area of research [46, 45, 44, 3, 35]. The main idea behind these approaches is to shift from a global image similarity to a local similarity. Local similarity can be enforced based on the location of pixels in an image [50, 44, 35], local pixel features [45], or both [29, 3, 37]. Local similarity can also be enforced in a region-based manner [19, 24, 49, 46]. These methods outperform those based solely on image-level global similarity in downstream dense tasks.
Encoder-Decoder pre-training: Encoder-decoder architectures with skip connections between the encoder and decoder layers (e.g., U-Net-like models) have achieved state-of-the-art results in many segmentation tasks [40, 27, 53]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [30, 4, 6]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder. Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder, even though a randomly initialized decoder is later attached to the pre-trained encoder in downstream dense tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [14, 38, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn a data representation.
Figure 1: A: DeCon-SL. Instead of a classical encoder-only pre-training, a decoder is pre-trained alongside the encoder. We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of the encoder and decoder losses. B: DeCon-ML. Instead of computing the decoder loss at a single level, it is calculated across multiple levels (four in this figure). Additionally, a channel-wise dropout is applied at the output of each encoder level before it is passed through the skip connection to the decoder.
3 Methods
We introduce two architectural adaptations, as depicted in  Fig. 1: (A) a single-level decoder loss (DeCon-SL) and (B) a multi-level decoder loss (DeCon-ML).
3.1 DeCon-SL
Given a vision SSL framework that pre-trains an encoder using a teacher-student architecture [25], we retain the encoders and any existing “auxiliary layers” (e.g., predictors, projection heads), and add decoders and their corresponding auxiliary layers for both teacher and student networks. For all models, the auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size. We obtain two losses, one is computed from the encoder features while the other one is computed from the decoder features. Consequently, the loss function is defined as follows:
$Loss=\alpha\times L_{enc}+(1-\alpha)\times L_{dec}$ (1)
where, $\alpha$ represents the contribution of the encoder loss in the objective function.
The DeCon-SL architecture is illustrated in Fig. 1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent teacher and student encoders, respectively. $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the teacher and student encoders, respectively. Auxiliary layers can include any projector, predictor heads, or any layers used in the loss definition of the original SSL framework. $g_{\theta}$ and $g_{\epsilon}$ are the decoders of the teacher and student networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ are the auxiliary layers feeding from their respective decoder. $g_{\theta}$ and $g_{\epsilon}$ take the output of encoder $f_{\theta}$ and $f_{\epsilon}$ as input, respectively. Auxiliary layers $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are responsible for generating a representation from the encoder output, that is used to compute a contrastive loss for the encoders $L^{encoder}$. The contrastive loss, $L^{decoder}$ is calculated from the representation obtained from the auxiliary layers of the decoders, i.e., $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$. $f_{\epsilon}$ and $g_{\epsilon}$ are updated using the exponential moving average (EMA) of $f_{\theta}$ and $g_{\theta}$, respectively.
3.2 DeCon-ML
To address the challenge posed by skip connections between the encoder and decoder layers (as mentioned in the Related works section), we introduce a multi-layer contrastive decoder loss, consisting of a single loss for the encoder and multiple loss components for the decoder along with channel dropout and deep supervision, which we describe in detail below. Figure 1 B illustrates the DeCon-ML adaptation of a given SSL framework.
Channel dropout: The dropout layer is typically used to prevent overfitting, improve generalization, and act as an implicit ensemble [41, 43]. However, we extend its use by applying dropout to channels transferred through skip connections between the encoder and decoder. This prevents the model from relying heavily on specific features shared through skip connections and enforces full encoder training, leading to richer data representations at the bottleneck layer of the encoder. In DeCon-ML, we apply channel dropout to the output layers of the encoder at the different levels when passing these outputs to the decoder. No channel is zeroed-out when the inputs pass through the encoder. This channel dropout promotes a more comprehensive utilization of the encoder parameters and enhances representation power at various encoder levels while preserving the input information content as it passes through the encoder to its bottleneck.
Decoder deep supervision: An encoder-decoder SSL framework allows us to retrieve meaningful features from multiple levels of the decoder. We can then pre-train the architecture with multiple losses using decoder deep supervision. If the decoder is connected to the encoder at multiple levels through lateral/skip connections, decoder deep supervision strengthens the representation power of the encoder at different levels [39]. Since a decoder passes the encoder bottleneck representation in a bottom-up manner to its upper level layers, the final encoder bottleneck representation remains a non-negligible objective of the training since it impacts all decoder level losses. In contrast, applying deep supervision at multiple encoder stages in an encoder-only SSL framework might prevent the creation of a strong semantic representation at the encoder bottleneck. To perform the decoder deep supervision, we add auxiliary layers to each decoder level, and a loss is computed at each level between the teacher and student representations. When performing decoder deep supervision, the encoder loss will remain the same as Eq. 1, but the decoder loss will be updated as follows:
$Loss=$ $\alpha\times L_{enc}+(1-\alpha)\times L_{dds}$ (2)
$L_{dds}=$ $\frac{1}{j}\times\sum_{i=1}^{j}L_{dec_{i}}$ (3)
where $j$ is the number of decoder levels at which a loss will be calculated and $L_{dec_{i}}$ is the loss computed at level $i$. The final decoder loss in Eq. 2 is thus an average of the losses at the different levels.
3.3 Implementation details
Architecture: Following common practice in contrastive SSL approaches literature, we use a ResNet-50 encoder [21] as a backbone, a Feature Pyramid Network (FPN) [32] for downstream detection and instance segmentation, and a Fully Convolutional Network (FCN) [33] for downstream semantic segmentation [45, 22, 46] tasks. We implement FCN and skip-connected FPN as two decoder architectures in the DeCon framework adaptation. We follow the mmseg [11] FCN implementation used in SlotCon [46] and customize FPN implementation from Detectron2 [47] version 0.6. We use DeCon to adapt the SlotCon [46] and DenseCL [45] frameworks. More implementation details are available in Supplementary Material Section 1.
For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged. Two decoders, a student and a teacher decoder (either FCN or FPN), that take the output of their corresponding encoder as input are added to the framework. The auxiliary layers are replicated at the decoder level. In the case of decoder deep supervision with FPN decoders, auxiliary layers are replicated independently at the four decoder levels. No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder. The decoder projectors differ from the encoder projectors only in input channels: 2048 for the encoder and 256 for the FPN and FCN decoders. We follow SlotCon by updating the teacher branch and the encoder and decoder slot predictors with Backpropagation, the student branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training. Our implementation is based on the official SlotCon repository. Detailed figures of SlotCon DeCon-SL and DeCon-ML adaptations are available in the Supplementary Material Section 1.
For DenseCL, the two encoders, along with their respective auxiliary layers—i.e., the global and dense projection heads—are kept the same as well as the independent dictionaries for global and dense losses. Additionally, we introduce two decoders, which take their respective encoders’ outputs as inputs. We also replicate the encoder’s dense and global projection heads and dictionaries separately for the two decoders. As for SlotCon, the input dimension of the decoder projectors is 256 instead of 2048 for the encoder. The encoders and decoders dictionaries size is set to 16384. We follow DenseCL by updating the first branch of the architecture through Backpropagation, while the second branch is updated using an EMA of the first one. Due to extensive memory requirements of replicating dictionaries at each decoder level, we do not conduct decoder deep supervision or channel dropout (DeCon-ML) experiments with DenseCL. We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0.
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31]. The augmentations and all the hyperparameters—learning rate, weight decay, schedulers, batch size, input shape, number of epochs, momentum, and so on—are kept the same as used in the original SSL framework: SlotCon or DenseCL. The SlotCon models are pre-trained for 800 epochs using a LARS optimizer, with a total batch size of 512 across all available GPUs. The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs. The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each). The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU. The DenseCL models are pre-trained for 800 epochs using an SGD optimizer with a total batch size of 256. The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum). The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each).
To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec. 3.4. For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with batch size. For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC. These pre-trainings are performed with one NVIDIA A6000 GPU with 48GB of memory.
3.4 Evaluation protocol
We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks. We either transfer the pre-trained encoder or both the encoder and decoder for each fine-tuning experiment. All fine-tuning results reported in this paper are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same dataset are run on a similar platform with the same hardware. We perform object detection and instance segmentation using Detectron2 package [47] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
Object detection and instance segmentation: We perform object detection on Pascal VOC dataset [18] and object detection and instance segmentation on COCO dataset [31]. We fine-tune a Faster R-CNN with a FPN C4 backbone using VOC trainval07+12 dataset for 24000 iterations and evaluate on the VOC test2007 set following previous works [22, 50]. For COCO object detection and instance segmentation, we fine-tune a Mask R-CNN with FPN backbone on the COCO train2017 dataset with 90000 iterations and evaluate on COCO val2017 split following SlotCon [46]. We initialize the network either with a pre-trained ResNet-50 encoder or with both a ResNet-50 encoder and a pre-trained FPN decoder, and fine-tune end-to-end in all cases. All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB memory.
Semantic segmentation on in-domain datasets: We perform semantic segmentation on Pascal VOC [18] and Cityscapes [13] datasets that contain scene-centric data following SlotCon [46] and PixCon [37]. A ResNet-50 encoder with a two-layer FCN decoder, similar to the FCN decoder used during pre-training, is used. For Pascal VOC, we train on VOC train_aug2012 set for 30000 iterations with a batch size of 16 and evaluate on VOC val2012 set. For Cityscapes, we fine tune on the train_fine set for 90000 iterations with a batch size of 16 and evaluate on the val_fine set. We initialize the network either with a pre-trained ResNet-50 encoder or with both the ResNet-50 encoder and the FCN decoder pre-trained and fine-tune end-to-end in all cases. All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB (Cityscapes) or 40GB (Pascal VOC) memory.
Semantic segmentation on out-of-domain datasets: We also evaluate generalizability to out-of-domain datasets for downstream tasks on different dataset sizes, considering 5, 25, and 100% randomly selected training samples. We fine-tune a ResNet-50 encoder coupled with an FCN decoder for semantic segmentation on REFUGE [36] and ISIC 2017 [10] dataset. REFUGE dataset contains 1200 retinal images in total representing 3 different classes: optic disk, optic cup, and background, divided into 400 training, 400 validation, and 400 testing samples. We train the ResNet-50-FCN network end-to-end for 80000 iterations with a batch size of 16 on the training set using one NVIDIA A100 GPU (40GB). The learning rate starts at 0.01 and decreases with a polynomial scheduler with a power of 0.9. We evaluate the model on the test set using the last iteration checkpoint and report the results.
ISIC 2017 is a skin-lesion segmentation dataset which consists of 2000 training, 150 validation, and 600 testing images. We fine-tune the same ResNet-50-FCN architecture on the training set for 24000 iterations on one NVIDIA V100 with 32GB of memory. Hyper parameters are the same as for REFUGE. We evaluate on the validation set every 500 iterations, select the best checkpoint and report the results of this checkpoint on the test set.
Table 1: Performance of DeCon-SL adaptation ($\alpha=0.5$) of SlotCon and DenseCL using FPN and FCN as options for decoders as well as baseline approaches. COCO dataset was used for pre-training. The same decoder architecture was used for pre-training and fine-tuning. The architecture was then fine-tuned on various segmentation and object detection tasks. All fine-tuning results were averaged over three runs. Up and down arrows indicate better and worst performance compared to baseline, respectively.
Framework Adaptation Decoder Loss Transfer Object Detection Object Detection Instance Seg. Semantic Seg.
VOC COCO COCO VOC City
$L_{enc}$ $L_{dec}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
SlotCon None None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
DeCon-SL FPN 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$ - -
55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$ - -
FCN - - - - - - - - - 72.42  $\uparrow$ 75.67  $\downarrow$
- - - - - - - - - 72.58  $\uparrow$ 76.00  $\uparrow$
DenseCL None None 54.50 80.47 59.97 39.60 59.30 43.29 35.65 56.49 38.25 71.11 75.87
DeCon-SL FPN 54.71  $\uparrow$ 80.65  $\uparrow$ 60.44  $\uparrow$ 39.60 59.36  $\uparrow$ 43.31  $\uparrow$ 35.74  $\uparrow$ 56.45  $\downarrow$ 38.40  $\uparrow$ - -
54.41  $\downarrow$ 80.24  $\downarrow$ 60.01  $\uparrow$ 39.55  $\downarrow$ 59.41  $\uparrow$ 43.15  $\downarrow$ 35.71  $\uparrow$ 56.44  $\downarrow$ 38.34  $\uparrow$ - -
FCN - - - - - - - - - 71.31  $\uparrow$ 76.04  $\uparrow$
- - - - - - - - - 71.59  $\uparrow$ 76.17  $\uparrow$
4 Results
DeCon-SL: Table 1 shows the results of DeCon-SL adaptation. DeCon pre-trained components outperformed baseline models in downstream object detection and segmentation when only the encoder was transferred, except for Cityscapes semantic segmentation with SlotCon. The transfer of both a pre-trained encoder and an FCN decoder always improved the performance further compared to transferring the pre-trained encoder only. Transferring the encoder and the FPN decoder pre-trained with DeCon-SL adaptation of DenseCL did not always result in improving model performance for downstream tasks.
DeCon-ML: Table 2 shows the results of DeCon-ML adaptation. We can see that adding the decoder deep supervision and a channel dropout probability of 0.5 during pre-training always enhanced fine-tuning performance compared to baseline approach (SlotCon) and DeCon-SL adaptation in all object detection and instance segmentation tasks.
Table 2: Performance of DeCon-ML adaptation of SlotCon with $\alpha=0.5$ and $dropout=0.5$ pre-trained on COCO dataset. We used an FPN as a DeCon decoder during pre-training and fine-tuning. We averaged all fine-tuning results over three runs. Arrows indicate performance gain or loss compared to baseline.
Frame work Adaptation Loss Transfer Object Detection Object Detection Instance Segmentation
VOC COCO COCO
$L_{enc}$ $L_{dec}$ $L_{dds}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
SlotCon None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
DeCon-SL 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$
55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$
DeCon-ML 55.60  $\uparrow$ 81.84  $\uparrow$ 61.83  $\uparrow$ 41.10  $\uparrow$ 61.37  $\uparrow$ 45.01  $\uparrow$ 37.05  $\uparrow$ 58.46  $\uparrow$ 39.87  $\uparrow$
55.57  $\uparrow$ 81.77  $\uparrow$ 61.76  $\uparrow$ 41.15  $\uparrow$ 61.38  $\uparrow$ 44.89  $\uparrow$ 37.05  $\uparrow$ 58.43  $\uparrow$ 39.88  $\uparrow$
Impact of the channel dropout: Impact of channel dropout is highlighted in Tab. 3. The decoder deep supervision on its own did not always improve downstream performance when only transferring the encoder in DeCon-ML. When transferring both encoder and decoder, downstream performance improved on Pascal VOC but not on COCO, which was the dataset used for pre-training. When the decoder deep supervision was coupled with channel dropout, performance improved compared to DeCon-SL. The channel dropout probability impacted the downstream performance differently when transferring only the encoder or both encoder and decoder. A channel dropout probability of 0.5 led to the best performance when transferring the encoder only. When transferring both encoder and decoder, a lower channel dropout probability of 0.25 led to the best performance on Pascal VOC object detection, and COCO instance segmentation. Comparing the performance of DeCon-ML adaptation with an encoder-only pre-training, we observed improvements in all scenarios regardless of dropout values.
Table 3: Impact of channel dropout probability on DeCon-ML. We pre-trained a DeCon-ML adaptation of SlotCon framework on COCO dataset with different channel dropout probabilities. We used an FPN as a DeCon decoder during pre-training and fine-tuning. We averaged all fine-tuning results over three runs.
Transfer SSL Loss $\alpha$ Dropout VOC Object Detection COCO Object Detection COCO Instance Seg.
$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
Enc SlotCon 1 0 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
DeCon-SL 0.5 0 55.57 81.75 61.76 40.94 60.97 44.82 36.92 58.03 39.80
0.5 55.20 81.56 61.33 40.90 61.23 44.64 36.91 58.30 39.67
DeCon-ML 0 55.52 81.67 62.01 40.85 60.95 44.67 36.83 58.12 39.55
0.25 55.54 81.88 61.10 40.95 61.13 44.88 36.97 58.22 39.76
0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87
0.75 55.29 81.64 61.36 41.02 61.27 44.96 37.05 58.33 39.85
Enc + Dec DeCon-SL 0.5 0 55.40 81.48 61.21 41.05 61.34 44.85 37.03 58.33 39.86
DeCon-ML 0 55.70 81.89 61.69 40.97 61.37 44.66 37.08 58.45 39.88
0.25 55.79 81.79 61.93 41.06 61.36 44.99 37.13 58.52 39.86
0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88
0.75 55.37 81.57 61.48 41.11 61.45 44.91 37.13 58.55 39.99
Impact of encoder loss weight in DeCon: As shown in Tab. 4, the encoder loss weight $\alpha$ played a crucial role in downstream performance. Decreasing $\alpha$ down to 0 for DeCon-ML pre-training and down to 0.25 for DeCon-SL pre-training increased the fine-tuning performance. This pattern was also observed when transferring both the encoder and the decoder. The downstream performance of DeCon adaptation was higher than that of encoder-only pre-training (baseline, $\alpha=1$) for object detection and instance segmentation tasks, whether transferring only the encoder or both the encoder and decoder. For Cityscapes semantic segmentation, no pattern was observed regarding the relationship between the value of $\alpha$ and model performance.
Table 4: Impact of the encoder weight $\alpha$. We pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder ($dropout=0.5$) and a DeCon-SL adaptation of SlotCon with an FCN decoder on COCO dataset. We averaged all fine-tuning results over three runs.
Transfer $\alpha$ SSL strategy Object Detection Object Detection Instance Seg. SSL strategy Semantic Seg.
VOC COCO COCO VOC City
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
Enc 1 SlotCon 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 SlotCon 71.50 75.95
0.75 DeCon-ML-FPN 0.5 Dropout ($L_{enc}+L_{dds}$) 55.33 81.68 61.34 40.98 61.28 44.76 36.95 58.34 39.72 DeCon-SL-FCN ($L_{enc}+L_{ds})$ 71.98 76.18
0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87 72.42 75.62
0.25 55.69 81.98 62.05 41.05 61.19 44.85 37.08 58.29 39.70 73.01 76.21
0 55.70 81.91 62.07 41.18 61.38 44.91 37.12 58.35 39.94 72.17 75.71
Enc + Dec 0.75 55.46 81.51 61.35 40.96 61.10 44.87 36.90 58.07 39.73 72.80 76.06
0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88 72.58 75.97
0.25 55.67 81.91 61.62 41.17 61.41 44.87 37.13 58.46 39.76 72.96 76.28
0 55.61 81.83 61.60 41.21 61.39 44.99 37.11 58.41 39.92 71.61 75.38
Heterogeneous pre-training and fine-tuning decoders: All the previous fine-tuning results were obtained using the pre-trained encoder of a DeCon adaptation, where the pre-training decoder matched the fine-tuning decoder. However, Tab. 5 shows that even if we pre-trained a DeCon adaptation of SlotCon with an FCN decoder and transferred the encoder only for an FPN-based downstream task, we obtained better performance than a baseline SlotCon pre-training. It is also true for the opposite scenario, when FPN decoder was used for pre-training DeCon-ML, and FCN decoder was used for fine-tuning.
Table 5: Heterogeneous pre-training and fine-tuning decoders. We pre-trained a DeCon-SL adaptation of SlotCon with an FCN decoder and transferred the pre-trained encoder to FPN-based models for downstream tasks. Similarly, we pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder and transferred the pre-trained encoder to FCN-based models for downstream tasks. We averaged fine-tuning results over three runs.
Pretraining $\alpha$ Drop Loss Obj. Det. Obj. Det. Inst. Seg. Semantic Seg.
VOC COCO COCO VOC City
$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP AP mIoU mIoU
SlotCon 1 0 55.16 40.81 36.80 71.50 75.95
DeCon-ML-FPN 0 0.5 - - - 72.92 76.45
DeCon-SL-FCN 0.5 0 55.29 40.96 36.98 - -
Comparison with state-of-the-art: Table 6 shows that our top-performing DeCon-ML and DeCon-SL adaptations achieved better performance than state-of-the-art methods on COCO object detection and instance segmentation, and VOC semantic segmentation tasks. A more detailed comparison is available in Supplementary Material Section 2.
[…]
Method Obj. Det. Inst. Seg. Semantic Seg.
VOC COCO COCO VOC City
APb APb APm mIoU mIoU
SlotCon ‡ [46] 55.16 40.81 36.80 71.50 75.95
PixCon-SR † [37] 57.55 40.81 36.84 72.95 76.62
DeCon-ML-SlotCon-FPN [ours] 55.70 41.18 37.12 72.92 76.45
DeCon-SL-SlotCon-FCN [ours] - - - 73.01 76.21
DeCon-SL performance in out-of-domain tasks: Table 7 depicts the performance of a model pre-trained on COCO to out-of-domain datasets. We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased. Further improvement was achieved when transferring the pre-trained decoder along with the encoder for ISIC dataset. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if we transferred only the encoder. We also observed that pre-trained models outperformed models trained with randomly initialized weights. This gap was more notable in REFUGE dataset (over 20%).
Table 7: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights. We pre-trained SlotCon and its DeCon-SL adaptation with an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets. We fine-tuned on REFUGE and ISIC datasets under various dataset sizes. We averaged the fine-tuning mIoU results over three runs.
SSL strategy Loss Transfer Pre-trained on COCO Pre-trained on Downstream Data
Refuge ISIC Refuge ISIC
$L_{enc}$ $L_{dec}$ enc dec 5% 25% 100% 5% 25% 100% 100% 100%
Random init. - - - - 49.53 41.41 62.82 75.66 78.38 80.65 62.82 80.65
SlotCon 69.60 77.80 82.83 74.95 79.05 82.52 79.86 83.19
DeCon-SL 71.75 78.92 83.57 75.62 79.56 82.84 80.71 83.66
72.09 77.27 83.25 76.00 79.97 82.94 81.01 83.25
5 Discussion
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation (Tab. 6). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab. 5). We showed that pre-training an FPN DeCon adaptation of SlotCon leads to better results when using deep supervision (DeCon-ML) compared to the baselines as well as DeCon-SL, both when transferring encoder only and when transferring encoder and decoder. Framework specific hyperparameter tuning of the auxiliary layers could increase further the performance, as shown in Supplementary material Section 3.
We introduced a combined loss where encoder and decoder loss contribute to the total loss in a weighted manner. We observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks. This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a construction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28]. Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. We also observed in Tab. 3 that dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML).
As shown in Tab. 1, DeCon-SL adaptation of an SSL framework also allows for pre-training a decoder efficiently. However, transferring the FPN decoder along the encoder for VOC object detection did not increase detection performance compared to transferring only the encoder, potentially indicating a form of over-fitting of the FPN decoder to the pre-training dataset. This is alleviated by deep supervision in DeCon-ML adaptation as shown in Tab. 3. We showed that DeCon can be used to adapt two different SSL frameworks using different concepts to learn local information (pixel level VS object/group level). This suggests that the proposed DeCon adaptation can be used in any SSL contrastive method designed for dense downstream tasks.
While DeCon-SL adaptation with an FCN decoder benefits DenseCL framework, the added value of the FPN decoder transfer in DeCon-SL adaptation of DenseCL was not obvious on object detection and instance segmentation downstream tasks. However, the added value of the FPN decoder transfer in DeCon-SL adaptation of SlotCon was present across all scenarios, enhancing model performance, possibly due to the region-based contrastive mechanism of SlotCon. We could not investigate the impact of DeCon-ML adaptation on DenseCL due to its extensive memory requirements, which can be attributed to its reliance on a large number of negative pairs.
DeCon’s performance on VOC object detection and Cityscapes semantic segmentation is constrained by its base SSL framework, SlotCon, preventing it from surpassing the state-of-the-art (PixCon) for these tasks(Tab. 6). As an adaptation, DeCon’s effectiveness largely depends on the underlying framework. While integrating DeCon with the latest state-of-the-art framework could set a new benchmark across all datasets, we were unable to implement PixCon due to the unavailability of its code [37].
DeCon adaptation of an SSL framework was also advantageous for fine-tuning on small scale datasets and out-of-domain dense tasks compared to the encoder-only baseline framework (see Tab. 7). This benefit was also observed when pre-training directly on the target dataset. In such cases, DeCon is valuable compared to an encoder-only pre-training whether we transfer the pre-trained decoder or not.
One limitation of our approach is its high memory and computational demand. Pre-training a decoder using DeCon can significantly increase the parameter count and training time, which can restrict this approach to certain SSL frameworks. For example, the DeCon-SL adaptation of SlotCon with FPN decoder could not fit in a single 80GB memory GPU. More details on the parameter count are available in Supplementary Material Section 4. Another limitation of our work is that we only considered ResNet-50 backbone; however, we expect achieving superior results when using more efficient backbones such as EfficientNet [42].
Brempong et al. [4] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed SSL adaptation to offer a multi-stage continual pre-training in a contrastive manner.
6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation. We showed that not-only does this adaptation provide a solution for pre-training decoders, but it also improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating these enhancements into an existing framework, we demonstrated that DeCon outperforms or matches state-of-the-art methods across multiple downstream dense prediction tasks.

### v2

@@
-Should we pre-train a decoder in contrastive learning for dense prediction tasks?
+Beyond the Encoder: Joint Encoder-Decoder Contrastive Pre-Training Improves Dense Prediction
 Abstract
-Contrastive learning in self-supervised settings primarily focuses on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks.
-This conventional approach, however, overlooks the potential benefits of jointly pre-training both the encoder and decoder.
-In this paper, we propose DeCon: a framework-agnostic adaptation to convert an encoder-only self-supervised learning (SSL) contrastive approach to an efficient encoder-decoder framework that can be pre-trained in a contrastive manner.
-We first update the existing architecture to accommodate a decoder and its respective contrastive loss.
-We then introduce a weighted encoder-decoder contrastive loss with non-competing objectives that facilitates the joint encoder-decoder architecture pre-training.
-We adapt two established contrastive SSL frameworks tailored for dense prediction tasks, achieve new state-of-the-art results in COCO object detection and instance segmentation, and match state-of-the-art performance on Pascal VOC semantic segmentation.
-We show that our approach allows for pre-training a decoder and enhances the representation power of the encoder and its performance in dense prediction tasks.
-This benefit holds across heterogeneous decoder architectures between pre-training and fine-tuning and persists in out-of-domain, limited-data scenarios.
+Contrastive learning methods in self-supervised settings have primarily focused on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks.
+However, this conventional approach overlooks the potential benefits of jointly pre-training both encoder and decoder.
+In this paper, we propose DeCon, an efficient encoder-decoder self-supervised learning (SSL) framework that supports joint contrastive pre-training.
+We first extend existing SSL architectures to accommodate diverse decoders and their corresponding contrastive losses.
+Then, we introduce a weighted encoder-decoder contrastive loss with non-competing objectives to enable the joint pre-training of encoder-decoder architectures.
+By adapting an established contrastive SSL framework for dense prediction tasks, DeCon achieves new state-of-the-art results: on COCO object detection and instance segmentation when pre-trained on COCO dataset; across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K.
+Our results demonstrate that joint pre-training enhances the representation power of the encoder and improves performance in dense prediction tasks.
+This gain persists across heterogeneous decoder architectures, various encoder architectures, and in out-of-domain limited-data scenarios.
 **footnotetext: Equal contributions for this work. Co-first authors are ordered in alphabetical order.
 1 Introduction
-The increasing demand for efficient Deep Learning (DL) methods arises from their ability to address complex tasks.
-While traditional approaches like supervised learning can yield powerful, high-performing models, they depend heavily on large volumes of high-quality labeled data.
-However, obtaining such annotations is often tedious, expensive, and sometimes impractical, even when abundant data is available.
-To overcome this bottleneck, one can benefit from a pre-trained model.
-Pre-trained models are usually models trained from scratch on a large scale dataset and allow for better performance on downstream tasks with smaller datasets.
+The growing demand for efficient Deep Learning (DL) methods stems from their ability to address complex tasks.
+While traditional approaches like supervised learning can yield high-performing models, they depend on large volumes of high-quality annotated data.
+Obtaining such annotations is often labor-intensive, expensive, and sometimes impractical, even when abundant data is available.
+A common workaround is to use pre-trained models, typically encoders trained from scratch on large-scale datasets, to boost performance on downstream tasks with limited labels.
 Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation.
-During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on part of or on a full annotated dataset.
-ImageNet [15] is a large and widely utilized dataset with annotated classes.
-However, models pre-trained on ImageNet classification are not the best fit for every downstream tasks, especially when the downstream objective is not classification.
-Since creating large-scale annotated dataset for every tasks presents numerous challenges, Self-Supervised Learning (SSL) has emerged as a promising alternative, enabling models to learn from the vast amount of unlabeled data and perform effectively on various downstream tasks.
-These pre-trained models can reach downstream performance close to a fully supervised approach using much smaller to no labeled datasets.
+During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on an annotated dataset.
+ImageNet [15] is a widely utilized labeled dataset for such pre-training.
+However, models pre-trained on ImageNet classification task often transfer poorly to downstream dense prediction tasks, such as detection and segmentation [56, 47, 41].
+Creating large-scale annotated datasets for every task is impractical.
+Self-Supervised Learning (SSL) is a promising alternative, enabling models to learn from the large-scale unannotated data to enhance performance on various downstream tasks.
+These pre-trained models can reach downstream performance close to a fully supervised approach while using significantly fewer labeled examples.
 Notable approaches such as SimCLR [7], VicReg [2], and MoCo [22] showed promising results by leveraging contrastive learning for downstream classification.
-However, most SSL frameworks primarily target classification downstream tasks.
-Some efforts such as PixCon [37] or DenseCL [45], focus on dense prediction tasks like segmentation or object detection.
-In all of these efforts, authors only train the encoders of the DL architectures and focus on incorporating a local loss with existing classification-targeted SSL approaches to adapt with dense tasks.
-In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, a framework agnostic encoder-decoder adaptation to pre-train an encoder jointly with a decoder, improving its representation power and efficiently preparing it to be combined with a decoder in downstream tasks.
-Our contributions are as follows.
+However, most SSL frameworks primarily target classification-based downstream tasks.
+Some efforts, such as PixCon [39] and DenseCL [47], focus on dense prediction tasks like segmentation or object detection.
+However, these approaches only pre-train encoders for downstream use and rely on local losses adapted from classification-oriented SSL methods without jointly training the decoder.
+In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, an encoder-decoder framework that jointly pre-trains encoder and decoder in an SSL manner.
+DeCon  enhances the representation power of the encoder and prepares it more effectively for integration with decoders in downstream dense prediction tasks.
+The contributions of this paper are as follows:
   • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks.
-We show that jointly pre-training the encoder with a decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even without transferring the pre-trained decoder.
-  • We improve on this adaptation by introducing DeCon-ML, which extends DeCon-SL with a multi-level decoder loss function and an encoder level-wise channel dropout to promote a comprehensive feature usage of the encoder.
-It achieves state-of-the-art results on COCO object detection and instance segmentation and matches top performance on Pascal VOC semantic segmentation.
-  • We demonstrate that DeCon improves the encoder’s downstream performance compared to encoder-only frameworks, even when pre-training and fine-tuning decoders differ.
-  • We show that DeCon’s efficiency persists in out-of-domain downstream tasks and provides even greater benefits in scenarios with limited annotated fine-tuning data.
+We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained decoder.
+  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive feature utilization of the encoder.
+DeCon-ML achieves new state-of-the-art results in COCO object detection and instance segmentation when pre-trained on COCO dataset; and across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K.
+  • We demonstrate that DeCon improves the encoder performance in downstream tasks compared to encoder-only frameworks, even when the decoders used for pre-training and fine-tuning differ.
+  • We show that DeCon can achieve gains without increasing parameter count or GPU requirements relative to the original framework.
+  • We demonstrate that DeCon remains effective in out-of-domain downstream tasks and yield even greater benefits in scenarios with limited annotated data for fine-tuning.
 2 Related Works
-Self-Supervised pre-training: In SSL, deep learning models are trained on pretext tasks where the data itself provides supervision, which allow models to learn meaningful representations without any manual annotations.
-Early SSL works focused on designing these tasks, such as predicting missing parts of an image (image inpainting) [38], context prediction [16], solving jigsaw puzzles [34], colorizing grayscale images [52], and predicting rotations [20].
-These methods aimed to create surrogate objectives that encourage models to develop feature representations, which could later be transferred to downstream tasks.
-However, these methods often produced task-specific representations with limited generalizability and restricted scalability.
-Recent methods prioritize generalizability and move away from hand-crafted pretext tasks.
+Self-Supervised pre-training: In early SSL works, deep learning models were trained on pretext tasks where supervision was derived directly from the data itself.
+This enabled the models to learn meaningful representations without relying on manual annotations.
+These tasks included predicting missing parts of an image [40], context prediction [16], solving jigsaw puzzles [36], colorizing grayscale images [54], and predicting rotations [20].
+These methods aimed to create surrogate objectives that encouraged models to develop feature representations transferable to downstream tasks.
+However, they often produced task-specific representations with limited generalizability.
+Recent SSL methods prioritize generalizability and move away from hand-crafted pretext tasks.
 These approaches can be broadly categorized into generative and contrastive methods.
-Generative approaches in SSL aim to learn meaningful representations by modeling the underlying data distribution.
-This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 51], or adversarial methods, where a generative model learns to synthesize realistic samples while an encoder captures high-level features [48, 26].
+Generative approaches aim to learn meaningful representations by modeling the underlying data distribution.
+This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 53], or through adversarial methods, where a generator model learns to synthesize realistic samples to challenge a discriminator, thereby encouraging the encoder to extract informative semantic features [50, 26].
 Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image.
-To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—negative pairs [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum contrast [22] or other architectural and optimization techniques are employed [5].
-Dense pre-training: Even though most of the aforementioned frameworks can be used for downstream segmentation or detection, these models are not designed specifically for pixel-level, i.e., dense downstream tasks.
-Consequently, developing SSL approaches that generalize well to dense downstream tasks, such as the segmentation task, is an active area of research [46, 45, 44, 3, 35].
+To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—strategies such as negative sampling [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum encoders [22], and other architectural and optimization techniques have been employed [5].
+Dense pre-training: Although most of the aforementioned frameworks can be used for downstream segmentation or detection, they are not designed specifically for pixel-level—i.e., dense prediction downstream tasks such as instance and semantic segmentation.
+Consequently, developing SSL approaches that generalize well to such downstream tasks remains an active area of research [48, 47, 46, 3, 37].
 The main idea behind these approaches is to shift from a global image similarity to a local similarity.
-Local similarity can be enforced based on the location of pixels in an image [50, 44, 35], local pixel features [45], or both [29, 3, 37].
+They aim to enforce local similarity based on the location of pixels in an image [52, 46, 37], local pixel features [47], or both [29, 3, 39].
 Local similarity can also be enforced in a region-based manner [19, 24, 51, 48].
-These methods outperform those based solely on image-level global similarity in downstream dense tasks.
-Encoder-Decoder pre-training: Encoder-decoder architectures with skip connections between the encoder and decoder layers (e.g., U-Net-like models) have achieved state-of-the-art results in many segmentation tasks [40, 27, 53].
+These methods outperform those relying solely on image-level global similarity when evaluated on dense prediction downstream tasks.
+Encoder-only versus encoder-decoder pre-training learning: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [43, 27, 55].
 A common approach for pre-training such models in a self-supervised manner is image reconstruction [30, 4, 6].
@@
 Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors.
-However, they do not include a decoder, even though a randomly initialized decoder is later attached to the pre-trained encoder in downstream dense tasks to form a complete architecture.
-While some generative SSL frameworks use a decoder for pre-training [14, 38, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn a data representation.
+However, they do not include a decoder.
+After pre-training, a randomly initialized decoder is attached to the pre-trained encoder in dense prediction downstream tasks to form a complete architecture.
+While some generative SSL frameworks use a decoder for pre-training [14, 40, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn informative data representations.
 Figure 1: A: DeCon-SL.
 Instead of a classical encoder-only pre-training, a decoder is pre-trained alongside the encoder.
-We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of the encoder and decoder losses.
+We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of encoder and decoder losses.
 B: DeCon-ML.
@@
 The DeCon-SL architecture is illustrated in Fig.
-1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent teacher and student encoders, respectively.
-$aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the teacher and student encoders, respectively.
+1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent student and teacher encoders, respectively.
+$aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the student and teacher encoders, respectively.
 Auxiliary layers can include any projector, predictor heads, or any layers used in the loss definition of the original SSL framework.
-$g_{\theta}$ and $g_{\epsilon}$ are the decoders of the teacher and student networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ are the auxiliary layers feeding from their respective decoder.
+$g_{\theta}$ and $g_{\epsilon}$ are the decoders of the student and teacher networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ their respective auxiliary layers.
 $g_{\theta}$ and $g_{\epsilon}$ take the output of encoder $f_{\theta}$ and $f_{\epsilon}$ as input, respectively.
@@
 3.2 DeCon-ML
-To address the challenge posed by skip connections between the encoder and decoder layers (as mentioned in the Related works section), we introduce a multi-layer contrastive decoder loss, consisting of a single loss for the encoder and multiple loss components for the decoder along with channel dropout and deep supervision, which we describe in detail below.
+To promote a comprehensive usage of the encoder’s parameters, we introduce a multi-layer contrastive decoder loss.
+This updated framework consists of a single loss for the encoder and multiple loss components for the decoder in a deep supervision manner along with channel dropout, which we describe in detail below.
 Figure 1 B illustrates the DeCon-ML adaptation of a given SSL framework.
@@
 However, we extend its use by applying dropout to channels transferred through skip connections between the encoder and decoder.
-This prevents the model from relying heavily on specific features shared through skip connections and enforces full encoder training, leading to richer data representations at the bottleneck layer of the encoder.
-In DeCon-ML, we apply channel dropout to the output layers of the encoder at the different levels when passing these outputs to the decoder.
+In DeCon-ML, we apply channel dropout to the output layers of the different encoder levels by zero-ing out entire channels of the feature maps when passing these outputs to the decoder.
 No channel is zeroed-out when the inputs pass through the encoder.
-This channel dropout promotes a more comprehensive utilization of the encoder parameters and enhances representation power at various encoder levels while preserving the input information content as it passes through the encoder to its bottleneck.
+This channel dropout prevents the model from over-relying on specific features shared through skip connections.
+It encourages a more comprehensive use of the encoder’s parameters at the different levels, leading to a richer and more powerful data representation learned, while still preserving the input information content as it passes through the encoder to its bottleneck.
 Decoder deep supervision: An encoder-decoder SSL framework allows us to retrieve meaningful features from multiple levels of the decoder.
@@
 3.3 Implementation details
-Architecture: Following common practice in contrastive SSL approaches literature, we use a ResNet-50 encoder [21] as a backbone, a Feature Pyramid Network (FPN) [32] for downstream detection and instance segmentation, and a Fully Convolutional Network (FCN) [33] for downstream semantic segmentation [45, 22, 46] tasks.
-We implement FCN and skip-connected FPN as two decoder architectures in the DeCon framework adaptation.
-We follow the mmseg [11] FCN implementation used in SlotCon [46] and customize FPN implementation from Detectron2 [47] version 0.6.
-We use DeCon to adapt the SlotCon [46] and DenseCL [45] frameworks.
-More implementation details are available in Supplementary Material Section 1.
+Architecture: We use a ResNet-50 encoder [21] as a backbone and implement a Fully Convolutional Network (FCN) [34] and a skip-connected Feature Pyramid Network (FPN) [32] as two decoder architectures in the DeCon framework.
+We follow the mmseg [11] FCN implementation used in SlotCon [48] and customize the FPN implementation from Detectron2 [49] version 0.6.
+Details on these decoder architectures are provided in Supplementary Material Section 1.
+We use DeCon to adapt the SlotCon [48] framework.
+Additional experiments are conducted using DenseCL [47] and PixPro [52] framework or a ConvNeXt-small [33] backbone.
 For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged.
@@
 The auxiliary layers are replicated at the decoder level.
-In the case of decoder deep supervision with FPN decoders, auxiliary layers are replicated independently at the four decoder levels.
-No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder.
+In the case of decoder deep supervision with FPN decoders (DeCon-ML), auxiliary layers are replicated independently at the four decoder levels.
+No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder (DeCon-SL).
 The decoder projectors differ from the encoder projectors only in input channels: 2048 for the encoder and 256 for the FPN and FCN decoders.
-We follow SlotCon by updating the teacher branch and the encoder and decoder slot predictors with Backpropagation, the student branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training.
+We follow SlotCon by updating the student branch and the corresponding encoder and decoder slot predictors with Backpropagation, the teacher branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training.
 Our implementation is based on the official SlotCon repository.
-Detailed figures of SlotCon DeCon-SL and DeCon-ML adaptations are available in the Supplementary Material Section 1.
-For DenseCL, the two encoders, along with their respective auxiliary layers—i.e., the global and dense projection heads—are kept the same as well as the independent dictionaries for global and dense losses.
-Additionally, we introduce two decoders, which take their respective encoders’ outputs as inputs.
-We also replicate the encoder’s dense and global projection heads and dictionaries separately for the two decoders.
-As for SlotCon, the input dimension of the decoder projectors is 256 instead of 2048 for the encoder.
-The encoders and decoders dictionaries size is set to 16384.
-We follow DenseCL by updating the first branch of the architecture through Backpropagation, while the second branch is updated using an EMA of the first one.
-Due to extensive memory requirements of replicating dictionaries at each decoder level, we do not conduct decoder deep supervision or channel dropout (DeCon-ML) experiments with DenseCL.
-We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0.
-Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31].
-The augmentations and all the hyperparameters—learning rate, weight decay, schedulers, batch size, input shape, number of epochs, momentum, and so on—are kept the same as used in the original SSL framework: SlotCon or DenseCL.
-The SlotCon models are pre-trained for 800 epochs using a LARS optimizer, with a total batch size of 512 across all available GPUs.
+We adapt in a similar way DenseCL and PixPro into a DeCon-SL framework using an FCN decoder.
+More details are given in Supplementary Material 2.
+We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0 and PixPro from the original implementation.
+Although the backbone pre-trained remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML architecture.
+To enable a fair comparison with the original SlotCon framework using a similar parameter budget, we implemented a reduced version of DeCon-ML that includes only the first two decoder levels originating from the encoder bottleneck.
+Additionally, the hidden dimension in the decoder projector was reduced from 4096 to 2048.
+This smaller version, which matches the parameter count of the original SlotCon framework (when $\alpha=0$), is referred to as DeCon-ML-S, while the full model with four decoder levels and unmodified configuration is denoted DeCon-ML-L.
+Detailed figures of DeCon-SL, DeCon-ML-S and DeCon-ML-L adaptations of SlotCon are available in the Supplementary Material Section 3.
+Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31], COCO+ (train2017 + unlabeled2017 datasets) and ImageNet-1K datasets [15].
+The augmentations and hyperparameters—learning rate, weight decay, schedulers, input shape, momentum, and so on—are kept the same as used in the original SSL framework.
+Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet.
+The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 across all available GPUs.
 The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs.
+The ConvNeXt-small backbone is pre-trained in SlotCon and DeCon-ML using an AdamW optimizer [35], a base learning rate of 0.002 and weight-decay 0.05.
+Remaining hyperparameters are kept the same as for the ResNet-50 implementation.
 The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each).
 The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU.
-The DenseCL models are pre-trained for 800 epochs using an SGD optimizer with a total batch size of 256.
+For DenseCL, DeCon-SL is pre-trained using an SGD optimizer with a total batch size of 256.
 The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum).
 The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each).
+For PixPro, DeCon-SL is pre-trained using a LARS optimizer for 800 epochs with a batch size of 512 and a learning rate of 1.0 with a weight decay of 1e-5.
+The pre-trainings are conducted with eight NVIDIA A100 GPUs (40GB each).
 To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec.
 3.4.
-For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with batch size.
+For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with the batch size.
 For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC.
@@
 We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks.
-We either transfer the pre-trained encoder or both the encoder and decoder for each fine-tuning experiment.
+Unless stated otherwise, we only transfer the pre-trained encoder for the fine-tuning experiments.
+We initialize the network with a pre-trained backbone and fine-tune end-to-end in all cases.
 All fine-tuning results reported in this paper are averaged over three independent runs.
@@
 We perform object detection and instance segmentation using Detectron2 package [49] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
-Object detection and instance segmentation: We perform object detection on Pascal VOC dataset [18] and object detection and instance segmentation on COCO dataset [31].
-We fine-tune a Faster R-CNN with a FPN C4 backbone using VOC trainval07+12 dataset for 24000 iterations and evaluate on the VOC test2007 set following previous works [22, 50].
-For COCO object detection and instance segmentation, we fine-tune a Mask R-CNN with FPN backbone on the COCO train2017 dataset with 90000 iterations and evaluate on COCO val2017 split following SlotCon [46].
-We initialize the network either with a pre-trained ResNet-50 encoder or with both a ResNet-50 encoder and a pre-trained FPN decoder, and fine-tune end-to-end in all cases.
-All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB memory.
-Semantic segmentation on in-domain datasets: We perform semantic segmentation on Pascal VOC [18] and Cityscapes [13] datasets that contain scene-centric data following SlotCon [46] and PixCon [37].
-A ResNet-50 encoder with a two-layer FCN decoder, similar to the FCN decoder used during pre-training, is used.
-For Pascal VOC, we train on VOC train_aug2012 set for 30000 iterations with a batch size of 16 and evaluate on VOC val2012 set.
-For Cityscapes, we fine tune on the train_fine set for 90000 iterations with a batch size of 16 and evaluate on the val_fine set.
-We initialize the network either with a pre-trained ResNet-50 encoder or with both the ResNet-50 encoder and the FCN decoder pre-trained and fine-tune end-to-end in all cases.
-All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB (Cityscapes) or 40GB (Pascal VOC) memory.
-Semantic segmentation on out-of-domain datasets: We also evaluate generalizability to out-of-domain datasets for downstream tasks on different dataset sizes, considering 5, 25, and 100% randomly selected training samples.
-We fine-tune a ResNet-50 encoder coupled with an FCN decoder for semantic segmentation on REFUGE [36] and ISIC 2017 [10] dataset.
-REFUGE dataset contains 1200 retinal images in total representing 3 different classes: optic disk, optic cup, and background, divided into 400 training, 400 validation, and 400 testing samples.
-We train the ResNet-50-FCN network end-to-end for 80000 iterations with a batch size of 16 on the training set using one NVIDIA A100 GPU (40GB).
-The learning rate starts at 0.01 and decreases with a polynomial scheduler with a power of 0.9.
-We evaluate the model on the test set using the last iteration checkpoint and report the results.
-ISIC 2017 is a skin-lesion segmentation dataset which consists of 2000 training, 150 validation, and 600 testing images.
-We fine-tune the same ResNet-50-FCN architecture on the training set for 24000 iterations on one NVIDIA V100 with 32GB of memory.
-Hyper parameters are the same as for REFUGE.
-We evaluate on the validation set every 500 iterations, select the best checkpoint and report the results of this checkpoint on the test set.
-Table 1: Performance of DeCon-SL adaptation ($\alpha=0.5$) of SlotCon and DenseCL using FPN and FCN as options for decoders as well as baseline approaches.
-COCO dataset was used for pre-training.
-The same decoder architecture was used for pre-training and fine-tuning.
-The architecture was then fine-tuned on various segmentation and object detection tasks.
+Object detection and instance segmentation: We perform object detection and instance segmentation on COCO dataset [31]. We fine-tune a Mask R-CNN with FPN backbone following SlotCon [48].
+Semantic segmentation: We evaluate the semantic segmentation performance of a pre-trained ResNet-50 encoder with an FCN decoder on Pascal VOC [18] and Cityscapes [13] following SlotCon [48] and PixCon [39].
+Semantic segmentation on out-of-domain datasets: We assess generalization on REFUGE [38] and ISIC 2017 [10] datasets using the same ResNet-50-FCN architecture, fine-tuned on 5, 25, and 100% of the training data.
+More details on the training setups and hyperparameters are provided in the Supplementary Material Section 2.
+4 Results
+4.1 Experiments
+Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML Small and Large ($\alpha=0$ and $dropout=0.5$) adaptations of SlotCon trained with a ResNet-50 backbone.
+The architecture was pre-trained for 800 epochs on COCO and COCO+ and for 200 epochs on ImageNet-1K, and then fine-tuned on various segmentation and object detection tasks.
+Only the pre-trained encoder was transferred for the fine-tuning.
+Results for PixCon pre-trained on ImageNet-1K are not reported, as the corresponding checkpoint or evaluation was not made available in the original publication.
 All fine-tuning results were averaged over three runs.
-Up and down arrows indicate better and worst performance compared to baseline, respectively.
-Framework Adaptation Decoder Loss Transfer Object Detection Object Detection Instance Seg. Semantic Seg.
-VOC COCO COCO VOC City
-$L_{enc}$ $L_{dec}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
-SlotCon None None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
-DeCon-SL FPN 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$ - -
-55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$ - -
-FCN - - - - - - - - - 72.42  $\uparrow$ 75.67  $\downarrow$
-- - - - - - - - - 72.58  $\uparrow$ 76.00  $\uparrow$
-DenseCL None None 54.50 80.47 59.97 39.60 59.30 43.29 35.65 56.49 38.25 71.11 75.87
-DeCon-SL FPN 54.71  $\uparrow$ 80.65  $\uparrow$ 60.44  $\uparrow$ 39.60 59.36  $\uparrow$ 43.31  $\uparrow$ 35.74  $\uparrow$ 56.45  $\downarrow$ 38.40  $\uparrow$ - -
-54.41  $\downarrow$ 80.24  $\downarrow$ 60.01  $\uparrow$ 39.55  $\downarrow$ 59.41  $\uparrow$ 43.15  $\downarrow$ 35.71  $\uparrow$ 56.44  $\downarrow$ 38.34  $\uparrow$ - -
-FCN - - - - - - - - - 71.31  $\uparrow$ 76.04  $\uparrow$
-- - - - - - - - - 71.59  $\uparrow$ 76.17  $\uparrow$
-4 Results
-DeCon-SL: Table 1 shows the results of DeCon-SL adaptation.
-DeCon pre-trained components outperformed baseline models in downstream object detection and segmentation when only the encoder was transferred, except for Cityscapes semantic segmentation with SlotCon.
-The transfer of both a pre-trained encoder and an FCN decoder always improved the performance further compared to transferring the pre-trained encoder only.
-Transferring the encoder and the FPN decoder pre-trained with DeCon-SL adaptation of DenseCL did not always result in improving model performance for downstream tasks.
-DeCon-ML: Table 2 shows the results of DeCon-ML adaptation.
-We can see that adding the decoder deep supervision and a channel dropout probability of 0.5 during pre-training always enhanced fine-tuning performance compared to baseline approach (SlotCon) and DeCon-SL adaptation in all object detection and instance segmentation tasks.
-Table 2: Performance of DeCon-ML adaptation of SlotCon with $\alpha=0.5$ and $dropout=0.5$ pre-trained on COCO dataset.
-We used an FPN as a DeCon decoder during pre-training and fine-tuning.
-We averaged all fine-tuning results over three runs.
-Arrows indicate performance gain or loss compared to baseline.
-Frame work Adaptation Loss Transfer Object Detection Object Detection Instance Segmentation
-VOC COCO COCO
-$L_{enc}$ $L_{dec}$ $L_{dds}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
-SlotCon None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
-DeCon-SL 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$
-55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$
-DeCon-ML 55.60  $\uparrow$ 81.84  $\uparrow$ 61.83  $\uparrow$ 41.10  $\uparrow$ 61.37  $\uparrow$ 45.01  $\uparrow$ 37.05  $\uparrow$ 58.46  $\uparrow$ 39.87  $\uparrow$
-55.57  $\uparrow$ 81.77  $\uparrow$ 61.76  $\uparrow$ 41.15  $\uparrow$ 61.38  $\uparrow$ 44.89  $\uparrow$ 37.05  $\uparrow$ 58.43  $\uparrow$ 39.88  $\uparrow$
-Impact of the channel dropout: Impact of channel dropout is highlighted in Tab.
-3.
-The decoder deep supervision on its own did not always improve downstream performance when only transferring the encoder in DeCon-ML.
-When transferring both encoder and decoder, downstream performance improved on Pascal VOC but not on COCO, which was the dataset used for pre-training.
-When the decoder deep supervision was coupled with channel dropout, performance improved compared to DeCon-SL.
-The channel dropout probability impacted the downstream performance differently when transferring only the encoder or both encoder and decoder.
-A channel dropout probability of 0.5 led to the best performance when transferring the encoder only.
-When transferring both encoder and decoder, a lower channel dropout probability of 0.25 led to the best performance on Pascal VOC object detection, and COCO instance segmentation.
-Comparing the performance of DeCon-ML adaptation with an encoder-only pre-training, we observed improvements in all scenarios regardless of dropout values.
-Table 3: Impact of channel dropout probability on DeCon-ML.
-We pre-trained a DeCon-ML adaptation of SlotCon framework on COCO dataset with different channel dropout probabilities.
-We used an FPN as a DeCon decoder during pre-training and fine-tuning.
-We averaged all fine-tuning results over three runs.
-Transfer SSL Loss $\alpha$ Dropout VOC Object Detection COCO Object Detection COCO Instance Seg.
-$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
-Enc SlotCon 1 0 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
-DeCon-SL 0.5 0 55.57 81.75 61.76 40.94 60.97 44.82 36.92 58.03 39.80
-0.5 55.20 81.56 61.33 40.90 61.23 44.64 36.91 58.30 39.67
-DeCon-ML 0 55.52 81.67 62.01 40.85 60.95 44.67 36.83 58.12 39.55
-0.25 55.54 81.88 61.10 40.95 61.13 44.88 36.97 58.22 39.76
-0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87
-0.75 55.29 81.64 61.36 41.02 61.27 44.96 37.05 58.33 39.85
-Enc + Dec DeCon-SL 0.5 0 55.40 81.48 61.21 41.05 61.34 44.85 37.03 58.33 39.86
-DeCon-ML 0 55.70 81.89 61.69 40.97 61.37 44.66 37.08 58.45 39.88
-0.25 55.79 81.79 61.93 41.06 61.36 44.99 37.13 58.52 39.86
-0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88
-0.75 55.37 81.57 61.48 41.11 61.45 44.91 37.13 58.55 39.99
-Impact of encoder loss weight in DeCon: As shown in Tab.
-4, the encoder loss weight $\alpha$ played a crucial role in downstream performance.
-Decreasing $\alpha$ down to 0 for DeCon-ML pre-training and down to 0.25 for DeCon-SL pre-training increased the fine-tuning performance.
-This pattern was also observed when transferring both the encoder and the decoder.
-The downstream performance of DeCon adaptation was higher than that of encoder-only pre-training (baseline, $\alpha=1$) for object detection and instance segmentation tasks, whether transferring only the encoder or both the encoder and decoder.
-For Cityscapes semantic segmentation, no pattern was observed regarding the relationship between the value of $\alpha$ and model performance.
-Table 4: Impact of the encoder weight $\alpha$.
-We pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder ($dropout=0.5$) and a DeCon-SL adaptation of SlotCon with an FCN decoder on COCO dataset.
-We averaged all fine-tuning results over three runs.
-Transfer $\alpha$ SSL strategy Object Detection Object Detection Instance Seg. SSL strategy Semantic Seg.
-VOC COCO COCO VOC City
-AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
-Enc 1 SlotCon 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 SlotCon 71.50 75.95
-0.75 DeCon-ML-FPN 0.5 Dropout ($L_{enc}+L_{dds}$) 55.33 81.68 61.34 40.98 61.28 44.76 36.95 58.34 39.72 DeCon-SL-FCN ($L_{enc}+L_{ds})$ 71.98 76.18
-0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87 72.42 75.62
-0.25 55.69 81.98 62.05 41.05 61.19 44.85 37.08 58.29 39.70 73.01 76.21
-0 55.70 81.91 62.07 41.18 61.38 44.91 37.12 58.35 39.94 72.17 75.71
-Enc + Dec 0.75 55.46 81.51 61.35 40.96 61.10 44.87 36.90 58.07 39.73 72.80 76.06
-0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88 72.58 75.97
-0.25 55.67 81.91 61.62 41.17 61.41 44.87 37.13 58.46 39.76 72.96 76.28
-0 55.61 81.83 61.60 41.21 61.39 44.99 37.11 58.41 39.92 71.61 75.38
-Heterogeneous pre-training and fine-tuning decoders: All the previous fine-tuning results were obtained using the pre-trained encoder of a DeCon adaptation, where the pre-training decoder matched the fine-tuning decoder.
-However, Tab.
-5 shows that even if we pre-trained a DeCon adaptation of SlotCon with an FCN decoder and transferred the encoder only for an FPN-based downstream task, we obtained better performance than a baseline SlotCon pre-training.
-It is also true for the opposite scenario, when FPN decoder was used for pre-training DeCon-ML, and FCN decoder was used for fine-tuning.
-Table 5: Heterogeneous pre-training and fine-tuning decoders.
-We pre-trained a DeCon-SL adaptation of SlotCon with an FCN decoder and transferred the pre-trained encoder to FPN-based models for downstream tasks.
-Similarly, we pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder and transferred the pre-trained encoder to FCN-based models for downstream tasks.
-We averaged fine-tuning results over three runs.
-Pretraining $\alpha$ Drop Loss Obj. Det. Obj. Det. Inst. Seg. Semantic Seg.
-VOC COCO COCO VOC City
-$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP AP mIoU mIoU
-SlotCon 1 0 55.16 40.81 36.80 71.50 75.95
-DeCon-ML-FPN 0 0.5 - - - 72.92 76.45
-DeCon-SL-FCN 0.5 0 55.29 40.96 36.98 - -
-Comparison with state-of-the-art: Table 6 shows that our top-performing DeCon-ML and DeCon-SL adaptations achieved better performance than state-of-the-art methods on COCO object detection and instance segmentation, and VOC semantic segmentation tasks.
-A more detailed comparison is available in Supplementary Material Section 2.
-[…]
-Method Obj. Det. Inst. Seg. Semantic Seg.
-VOC COCO COCO VOC City
-APb APb APm mIoU mIoU
-SlotCon ‡ [46] 55.16 40.81 36.80 71.50 75.95
-PixCon-SR † [37] 57.55 40.81 36.84 72.95 76.62
-DeCon-ML-SlotCon-FPN [ours] 55.70 41.18 37.12 72.92 76.45
-DeCon-SL-SlotCon-FCN [ours] - - - 73.01 76.21
-DeCon-SL performance in out-of-domain tasks: Table 7 depicts the performance of a model pre-trained on COCO to out-of-domain datasets.
+means that we reimplemented the framework, * means that the results are taken from the original paper.
+Pret. Dataset Framework Pret. Dec. Object Detection Instance Segmentation Semantic Segmentation
+COCO COCO VOC City
+AP AP50 AP75 AP AP50 AP75 mIoU mIoU
+COCO PixCon [39]* - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62
+SlotCon [48] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
+DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21
+DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21
+DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45
+COCO+ PixCon [39]* - 41.2 - - 37.1 - - 73.9 77.0
+SlotCon [48] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43
+DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65
+DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00
+ImageNet-1K SlotCon [48] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15
+DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51
+ResNet-50 backbone: Table 1 presents the results of fine-tuning a ResNet-50 backbone pre-trained with the two best pre-training frameworks for dense downstream tasks and DeCon.
+A more comprehensive comparison with additional previously published methods is provided in the Supplementary Material 4.
+DeCon-ML-L is consistently establishing a new SATO for all tasks, except when pre-training is performed on COCO dataset for Pascal VOC and Cityscapes semantic segmentation fine-tuning, where it reduces the gap between PixCon and SlotCon.
+This supports the fact that the DeCon adpatation is robust across datasets.
+Secondly, we can see that all of DeCon versions consistently improve upon the base SlotCon framework for any of the pre-training datasets used.
+Notably, DeCon-ML-S, whose parameter count is similar to SlotCon, exhibits better performance than SlotCon.
+Table 2: Performance of DeCon-ML-L ($\alpha=0$ and $dropout=0.5$) adaptation of SlotCon using a ConvNeXt-small encoder.
+The architecture was pre-trained for 800 epochs on COCO, and then fine-tuned on various segmentation and object detection tasks.
+All fine-tuning results were averaged over three runs.
+Pret. Dataset Framework Pret. Dec. Obj. Det. Inst. Seg. Sem. Seg.
+COCO COCO VOC
+AP AP mIoU
+COCO SlotCon - 44.07 39.67 73.24
+DeCon-ML-L FPN 44.71 40.37 73.81
+ConvNeXt backbone: Table 2 presents the results of the fine-tuning of models pre-trained with a ConvNeXt-small backbone using DeCon-ML-L and the original SlotCon.
+A consistent improvement is seen with our encoder-decoder framework, confirming the robustness of our method to more modern and bigger backbones.
+Notably, the performance of the model as well as the performance improvement appear larger than for a smaller ResNet-50 backbone.
+Table 3: Performance of DeCon-SL ($\alpha=0.5$) adaptation of DenseCL and PixPro using a ResNet-50 encoder.
+The architectures were pre-trained for 800 epochs on COCO, and then fine-tuned on Pascal VOC and Cityscapes semantic segmentation.
+All fine-tuning results were averaged over three runs.
+Pret. Dataset Framework Pret. Dec. Transfer Sem. Seg.
+VOC City
+mIoU mIoU
+COCO DenseCL - Enc 71.11 75.87
+DeCon-SL (DenseCL) FCN Enc 71.31 76.04
+Enc + Dec 71.59 76.17
+PixPro - Enc 72.13 75.84
+DeCon-SL-(PixPro) FCN Enc 72.18 76.16
+Enc + Dec 73.08 76.19
+Generalizability across frameworks: Table 3 presents semantic segmentation fine-tuning results of models pre-trained with DenseCL and PixPro, and their DeCon-SL adaptation.
+The performance of the model pre-trained with a decoder is improved compared to the base framework and improves even further when the pre-trained decoder is transferred, supporting the hypothesis that DeCon could be adapted to multiple frameworks.
+Transferring the pre-trained decoder however, is not always beneficial to downstream performance in the case of SlotCon.
+More details can be found in Supplementary Material 5.
+Out-of-domain tasks: Table 4 depicts the fine-tuning performance of a model pre-trained on COCO to out-of-domain datasets.
 We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased.
@@
 However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training.
-In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if we transferred only the encoder.
+In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if only the encoder was transferred.
 We also observed that pre-trained models outperformed models trained with randomly initialized weights.
@@
 Table 4: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights.
-We pre-trained SlotCon and its DeCon-SL adaptation with an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets.
+We pre-trained SlotCon and its DeCon-SL adaptation with a ResNet-50 encoder and an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets.
 We fine-tuned on REFUGE and ISIC datasets under various dataset sizes.
@@
 72.09 77.27 83.25 76.00 79.97 82.94 81.01 83.25
+4.2 Ablation studies
+Table 5 presents an ablation study of the components in our DeCon-ML-L framework.
+The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, even when only the encoder is transferred, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision.
+This combination yields the most substantial gains, highlighting the critical role of channel dropout when paired with deep supervision in our approach.
+Table 5: Ablation of the different components of DeCon-ML-L and their impact on the downstream performance.
+We averaged all fine-tuning results over three runs.
+$\Delta$ is the performance gap between the previous framework and the framework updated with the corresponding new component.
+Framework Improvement COCO Obj Det. COCO Inst. Seg.
+AP $\Delta$ AP $\Delta$
+SlotCon 40.81 - 36.80 -
++FPN Decoder (Multi Level Loss) 40.85 +0.04 36.83 +0.03
++Dropout = 0.5 41.10 +0.25 37.05 +0.22
++Enc-Dec Loss Weight ($\alpha$) Tuning (DeCon-ML-L) 41.18 +0.08 37.12 +0.07
+Table 6 shows DeCon-ML performance with different values of encoder loss weight $\alpha$ and channel dropout.
+A channel dropout probability of 0.5 provided the best downstream performance for DeCon-ML-L.
+Hence, this value was chosen to ablate $\alpha$.
+An $\alpha$ value of 0 gave the best downstream performance meaning that the decoder loss can replace the encoder one and pre-train efficiently an encoder on its own.
+For DeCon-SL, an $\alpha$ value of 0.25 resulted in the best downstream performance, suggesting that having an encoder loss at the bottleneck is still beneficial to the pre-training in the absence of skip connections.
+Table 6: Impact of the channel dropout probability and of the encoder loss weight $\alpha$ on DeCon.
+We pre-trained DeCon adaptations of SlotCon framework on COCO dataset using different parameters and evaluate the pre-trained encoder fine-tuning on downstream tasks that use the same decoder architecture.
+We averaged all fine-tuning results over three runs.
+Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Frame work Deco der $\alpha$ Sem. Seg.
+COCO COCO VOC City
+AP AP mIoU mIoU
+Dropout
+DeCon-ML-L
+FPN
+0.50 0 40.85 36.83 - - - - -
+0.25 40.95 36.97 - - - - -
+0.50 41.10 37.05 - - - - -
+0.75 41.02 37.05 - - - - -
+$\alpha$
+0.75 0.50 40.98 36.90
+DeCon-SL (No Drop)
+FCN
+0.75 71.98 76.18
+0.50 41.10 37.05 0.50 72.42 75.62
+0.25 41.05 37.13 0.25 73.01 76.21
+0 41.18 37.11 0 72.17 75.71
 5 Discussion
+Table 7: COCO and COCO+ pretraining downstream performance improvement with DeCon-ML-L and Cohen’s d value. PixCon performance was extracted from the original paper.
+Method COCO obj det. COCO Inst. Seg. VOC Sem. Seg. City Sem. Seg.
+$\Delta$AP $\Delta$AP $\Delta$mIoU $\Delta$mIoU
+COCO Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
+SlotCon (2022) (Comp. to prev. works) +0.50 +0.40 -0.40 +0.10
+PixCon (2024) (Comp. to SlotCon) 0.00 +0.06 +1.30 +0.51
+DeCon-ML-L (Comp. to SlotCon) +0.37 +0.32 +1.42 +0.50
+Cohen’s d (DeCon-ML-L VS SlotCon) 2.35 1.94 6.37 2.51
+COCO+ Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
+SlotCon (2022) (Comp. to prev. works) +0.60 +0.50 Not avail. Not avail.
+PixCon (2024) (Comp. to SlotCon) -0.50 -0.50 -0.20 +0.40
+DeCon-ML-L (Comp. to SlotCon) +0.45 +0.27 +1.43 +0.57
+Cohen’s d (DeCon-ML-L VS SlotCon) 5.42 3.21 5.25 1.22
 In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework.
-We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation (Tab.
-6).
+We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation when pre-trained on COCO dataset, and consistently outperforms previous methods when pre-trained on COCO+, and ImageNet except for Cityscapes semantic segmentation (Tab.
+1).
 This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab.
-5).
-We showed that pre-training an FPN DeCon adaptation of SlotCon leads to better results when using deep supervision (DeCon-ML) compared to the baselines as well as DeCon-SL, both when transferring encoder only and when transferring encoder and decoder.
-Framework specific hyperparameter tuning of the auxiliary layers could increase further the performance, as shown in Supplementary material Section 3.
-We introduced a combined loss where encoder and decoder loss contribute to the total loss in a weighted manner.
-We observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks.
-This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a construction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28].
+1).
+We introduced a combined loss where encoder and decoder losses contribute to the total loss in a weighted manner.
+Figure 2: Pre-training loss dynamics for DeCon-ML-L and SlotCon.
+Figure 2 illustrates the loss behavior across different pre-trainings when both the encoder and decoder are used.
+As shown, decoder pre-training does not influence the encoder’s loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing.
+We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks.
+This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28].
+Additionally, Fig.
+3 presents the slots learned by SlotCon encoder and by the different decoder levels of DeCon-ML.
+It demonstrates that the decoder learns features that capture similar concepts as the encoder’s, while being more spatially precise.
+Figure 3: Slots, as defined in SlotCon, learned at different outputs of our architecture.
+The top left image from COCO validation dataset is overlayed with the slot that was the most represented in original SlotCon’s encoder output feature map, resized to the image’s shape.
+Other slots displayed are the slots that had the biggest overlap with SlotCon’s main encoder output slot.
+More details on the slots creation are available in Supplementary Material 6.
 Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder.
 The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training.
-We also observed in Tab.
-3 that dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML.
+In Supplementary Material Section 7, we provide experiments on using DeCon-SL with an FPN decoder.
+We observed that with an FPN decoder, dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML.
 Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML).
-As shown in Tab.
-1, DeCon-SL adaptation of an SSL framework also allows for pre-training a decoder efficiently.
-However, transferring the FPN decoder along the encoder for VOC object detection did not increase detection performance compared to transferring only the encoder, potentially indicating a form of over-fitting of the FPN decoder to the pre-training dataset.
-This is alleviated by deep supervision in DeCon-ML adaptation as shown in Tab.
-3.
-We showed that DeCon can be used to adapt two different SSL frameworks using different concepts to learn local information (pixel level VS object/group level).
-This suggests that the proposed DeCon adaptation can be used in any SSL contrastive method designed for dense downstream tasks.
-While DeCon-SL adaptation with an FCN decoder benefits DenseCL framework, the added value of the FPN decoder transfer in DeCon-SL adaptation of DenseCL was not obvious on object detection and instance segmentation downstream tasks.
-However, the added value of the FPN decoder transfer in DeCon-SL adaptation of SlotCon was present across all scenarios, enhancing model performance, possibly due to the region-based contrastive mechanism of SlotCon.
-We could not investigate the impact of DeCon-ML adaptation on DenseCL due to its extensive memory requirements, which can be attributed to its reliance on a large number of negative pairs.
-DeCon’s performance on VOC object detection and Cityscapes semantic segmentation is constrained by its base SSL framework, SlotCon, preventing it from surpassing the state-of-the-art (PixCon) for these tasks(Tab.
-6).
-As an adaptation, DeCon’s effectiveness largely depends on the underlying framework.
-While integrating DeCon with the latest state-of-the-art framework could set a new benchmark across all datasets, we were unable to implement PixCon due to the unavailability of its code [37].
-DeCon adaptation of an SSL framework was also advantageous for fine-tuning on small scale datasets and out-of-domain dense tasks compared to the encoder-only baseline framework (see Tab.
-7).
-This benefit was also observed when pre-training directly on the target dataset.
-In such cases, DeCon is valuable compared to an encoder-only pre-training whether we transfer the pre-trained decoder or not.
+Framework specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (shown in Supplementary Material Section 8).
+We demonstrated that our method is effective with a different type of encoders.
+Specifically, pre-training a ConvNeXt-small backbone, which is both more efficient and larger than the ResNet-50 backbone, also resulted in superior downstream performance when integrated with a decoder in the DeCon-ML-L framework.
+These results attest to the scalability and practical utility of the proposed approach for pre-training modern backbone architectures.
+DeCon was also used to successfully adapt three different SSL frameworks using different concepts (pixel level VS object/group level) to learn local information (Tab.
+3).
+This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods designed for dense downstream tasks.
+Table 8: Pre-training parameter count, GPU setup, and time required per epoch for the different architectures.
+We study various pre-training setups on the COCO dataset using the SlotCon framework.
+Module sizes are given in millions of parameters.
+Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU setup Time per epoch (s)
+SlotCon 23.51 x 2 0 71.51 0 118.52 1 * A100 80GB 255
+DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1 * A100 80GB 300
+DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1 * A100 80GB 280
+DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 8 * H100 80GB 165
+DeCon adaptation of an SSL framework provided notable advantages for fine-tuning on small-scale datasets and out-of-domain dense prediction tasks, outperforming the encoder-only baseline (see Tab.
+4).
+This improvement persisted even when pre-training was conducted directly on the small target datasets.
+Furthermore, DeCon achieved superior performance compared to related frameworks when only a limited fraction (5–25%) of the original labeled data was available for fine-tuning.
+These results highlight the value of DeCon in settings with very limited annotated samples or small-scale, domain-specific datasets, such as those in medical or agricultural applications, regardless of whether the pre-trained decoder is transferred.
+Performance of deep learning models can vary across multiple runs due to inherent stochasticity.
+To mitigate the impact of randomness while considering computational resource constraints, each experiment was conducted three times, with all fine-tunings for a specific downstream task performed on similar hardware.
+The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8> 0.8) in Tab.
+7, exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests.
+Details on the computation of these values are provided in Supplementary Material 10.
+Furthermore, the performance gains shown in Tab.
+7 are comparable to those reported in prior studies.
 One limitation of our approach is its high memory and computational demand.
-Pre-training a decoder using DeCon can significantly increase the parameter count and training time, which can restrict this approach to certain SSL frameworks.
-For example, the DeCon-SL adaptation of SlotCon with FPN decoder could not fit in a single 80GB memory GPU.
-More details on the parameter count are available in Supplementary Material Section 4.
-Another limitation of our work is that we only considered ResNet-50 backbone; however, we expect achieving superior results when using more efficient backbones such as EfficientNet [42].
+However, we have shown in Tab.
+1 that if GPU is a constraint, the user can decide to reduce the number of decoder levels used in the pre-training and tune the decoder projector layers to reduce the parameter count of DeCon-ML architecture and match the parameter count of the original framework (as shown in Tab.
+8), while still improving the performance of the pre-training.
 Brempong et al.
@@
 It could also result in suboptimal feature alignment and increased training complexity.
-Future works could involve extending our proposed SSL adaptation to offer a multi-stage continual pre-training in a contrastive manner.
+Future works could involve extending our proposed DeCon adaptation to offer a multi-stage continual pre-training in a contrastive manner.
 6 Conclusion
 In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation.
-We showed that not-only does this adaptation provide a solution for pre-training decoders, but it also improves the encoder’s representation significantly.
+We showed that this adaptation improves the encoder’s representation significantly.
 We proposed two variants of DeCon: DeCon-SL and DeCon-ML.
 The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power.
-By integrating these enhancements into an existing framework, we demonstrated that DeCon outperforms or matches state-of-the-art methods across multiple downstream dense prediction tasks.
+By integrating the proposed contrastive joint encoder-decoder pre-training strategy into an existing framework, we show that DeCon consistently outperforms or matches state-of-the-art methods across diverse dense prediction tasks, highlighting the effectiveness of this unified pre-training approach.
+Acknowledgments
+Computing resources were provided by the University of British Columbia Cloud Innovation Centre, powered by Amazon Web Services https://cic.ubc.ca/.
+This research was also partly enabled by support provided by the Digital Research Alliance of Canada https://alliancecan.ca/, and the Research Computing Services group at the University of Calgary https://rcs.ucalgary.ca/RCS_Home_Page.

### v3

@@
 Then, we introduce a weighted encoder-decoder contrastive loss with non-competing objectives to enable the joint pre-training of encoder-decoder architectures.
-By adapting an established contrastive SSL framework for dense prediction tasks, DeCon achieves new state-of-the-art results: on COCO object detection and instance segmentation when pre-trained on COCO dataset; across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K.
-Our results demonstrate that joint pre-training enhances the representation power of the encoder and improves performance in dense prediction tasks.
-This gain persists across heterogeneous decoder architectures, various encoder architectures, and in out-of-domain limited-data scenarios.
-**footnotetext: Equal contributions for this work. Co-first authors are ordered in alphabetical order.
+By adapting a contrastive SSL framework for dense prediction, DeCon establishes consistent state-of-the-art performance on most of the evaluated tasks when pre-trained on Imagenet-1K, COCO and COCO+.
+Notably, when pre-training a ResNet-50 encoder on COCO dataset, DeCon improves COCO object detection and instance segmentation compared to the baseline framework by +0.37 AP and +0.32 AP, respectively, and boosts semantic segmentation by +1.42 mIoU on Pascal VOC and by +0.50 mIoU on Cityscapes.
+These improvements generalize across recent backbones, decoders, datasets, and dense tasks beyond segmentation and object detection, and persist in out-of-domain scenarios, including limited-data settings, demonstrating that joint pre-training significantly enhances representation quality for dense prediction.
+Code is available at https://github.com/sebquetin/DeCon.git.
+**footnotetext: Equal contributions. Co-first authors are ordered in alphabetical order. Corresponding author: T. Ghosh (tapotosh.ghosh@ucalgary.ca)
 1 Introduction
 The growing demand for efficient Deep Learning (DL) methods stems from their ability to address complex tasks.
-While traditional approaches like supervised learning can yield high-performing models, they depend on large volumes of high-quality annotated data.
+While traditional approaches like supervised learning can train high-performing models, they depend on large volumes of high-quality annotated data.
 Obtaining such annotations is often labor-intensive, expensive, and sometimes impractical, even when abundant data is available.
@@
 In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, an encoder-decoder framework that jointly pre-trains encoder and decoder in an SSL manner.
-DeCon  enhances the representation power of the encoder and prepares it more effectively for integration with decoders in downstream dense prediction tasks.
+DeCon enhances the representation power of the encoder, and prepares it more effectively for integration with decoders in downstream dense prediction tasks.
 The contributions of this paper are as follows:
   • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks.
-We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained decoder.
-  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive feature utilization of the encoder.
-DeCon-ML achieves new state-of-the-art results in COCO object detection and instance segmentation when pre-trained on COCO dataset; and across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K.
-  • We demonstrate that DeCon improves the encoder performance in downstream tasks compared to encoder-only frameworks, even when the decoders used for pre-training and fine-tuning differ.
-  • We show that DeCon can achieve gains without increasing parameter count or GPU requirements relative to the original framework.
-  • We demonstrate that DeCon remains effective in out-of-domain downstream tasks and yield even greater benefits in scenarios with limited annotated data for fine-tuning.
-2 Related Works
+We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained encoder.
+  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive encoder feature utilization.
+  • DeCon achieves new state-of-the-art (SOTA) results in COCO object detection and instance segmentation, and Pascal VOC, Cityscapes and ADE20K semantic segmentation when pre-trained on ImageNet-1K and COCO+.
+  • We show that DeCon consistently improves performance across different backbones, a range of dense downstream tasks beyond object detection, semantic and instance segmentation as well as in various out-of-domain scenarios.
+  • We show that DeCon can achieve gains without increasing parameter count and with comparable GPU cost relative to the original framework, e.g. SlotCon [55].
+2 Related Work
 Self-Supervised pre-training: In early SSL works, deep learning models were trained on pretext tasks where supervision was derived directly from the data itself.
@@
 These approaches can be broadly categorized into generative and contrastive methods.
-Generative approaches aim to learn meaningful representations by modeling the underlying data distribution.
+Generative approaches learn representations by modeling the underlying data distribution.
 This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [24, 1, 20, 61], or through adversarial methods, where a generator model learns to synthesize realistic samples to challenge a discriminator, thereby encouraging the encoder to extract informative semantic features [57, 29].
@@
 These methods outperform those relying solely on image-level global similarity when evaluated on dense prediction downstream tasks.
-Encoder-only versus encoder-decoder pre-training learning: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [43, 27, 55].
+Encoder-only versus encoder-decoder pre-training: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [48, 30, 64].
 A common approach for pre-training such models in a self-supervised manner is image reconstruction [33, 5, 8].
@@
 After pre-training, a randomly initialized decoder is attached to the pre-trained encoder in dense prediction downstream tasks to form a complete architecture.
-While some generative SSL frameworks use a decoder for pre-training [14, 40, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn informative data representations.
+While some generative SSL frameworks use a decoder for pre-training [17, 43, 24], a clear gap remains in exploring vision model pre-training that jointly employs encoder and decoder losses within a unified contrastive learning framework to learn informative data representations.
 Figure 1: A: DeCon-SL.
@@
 Given a vision SSL framework that pre-trains an encoder using a teacher-student architecture [28], we retain the encoders and any existing “auxiliary layers” (e.g., predictors, projection heads), and add decoders and their corresponding auxiliary layers for both teacher and student networks.
-For all models, the auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size.
+For all models, the decoder auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size.
 We obtain two losses, one is computed from the encoder features while the other one is computed from the decoder features.
@@
 Architecture: We use a ResNet-50 encoder [26] as a backbone and implement a Fully Convolutional Network (FCN) [37] and a skip-connected Feature Pyramid Network (FPN) [34] as two decoder architectures in the DeCon framework.
-We follow the mmseg [11] FCN implementation used in SlotCon [48] and customize the FPN implementation from Detectron2 [49] version 0.6.
-Details on these decoder architectures are provided in Supplementary Material Section 1.
+We follow the mmseg [14] FCN implementation from SlotCon [55] and customize the Detectron2 [56] FPN implementation.
+Details on these decoders are provided in Supplementary Material Section 1.
 We use DeCon to adapt the SlotCon [55] framework.
-Additional experiments are conducted using DenseCL [47] and PixPro [52] framework or a ConvNeXt-small [33] backbone.
+Additional experiments are conducted using DenseCL [53] and PixPro [60] framework, or a ConvNeXt-Small [36] backbone.
 For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged.
@@
 We adapt in a similar way DenseCL and PixPro into a DeCon-SL framework using an FCN decoder.
-More details are given in Supplementary Material 2.
+More details are given in Supplementary Material Section 2.
 We adapt DenseCL implementation from mmselfsup package [15] version 1.0.0 and PixPro from the original implementation.
-Although the backbone pre-trained remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML architecture.
+Although the pre-trained backbone remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML pre-training.
 To enable a fair comparison with the original SlotCon framework using a similar parameter budget, we implemented a reduced version of DeCon-ML that includes only the first two decoder levels originating from the encoder bottleneck.
@@
 Pre-training setup: We pre-train DeCon on COCO 2017 dataset [35], COCO+ (train2017 + unlabeled2017 datasets) and ImageNet-1K datasets [18].
-The augmentations and hyperparameters—learning rate, weight decay, schedulers, input shape, momentum, and so on—are kept the same as used in the original SSL framework.
-Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet.
-The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 across all available GPUs.
-The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs.
-The ConvNeXt-small backbone is pre-trained in SlotCon and DeCon-ML using an AdamW optimizer [35], a base learning rate of 0.002 and weight-decay 0.05.
+The augmentations and hyperparameters—learning rate (LR), weight decay, schedulers, input shape, momentum, and so on—are kept the same as the adapted SSL framework.
+Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet-1K.
+The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 split across all available GPUs.
+The base LR is 1.0 scaled with the batch size following [55] and updated with a cosine LR decay schedule using a weight decay of $10^{-5}$ and 5 warm-up epochs.
+The ConvNeXt-Small backbone is pre-trained on COCO dataset with SlotCon and DeCon-ML using an AdamW optimizer [38], a base LR of 0.002 and weight-decay 0.05.
+To compare with ViT-based methods, we also pre-train a ConvNeXt-S backbone on ImageNet-1K for 250 epochs with DeCon-SL framework and an FCN decoder using a base LR of 0.001 and drop path rate of 0.1.
 Remaining hyperparameters are kept the same as for the ResNet-50 implementation.
-The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each).
+The pre-trainings using ConvNeXt-S backbones or FPN decoders are done using eight 80GB NVIDIA H100 GPUs.
 The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU.
-For DenseCL, DeCon-SL is pre-trained using an SGD optimizer with a total batch size of 256.
-The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum).
-The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each).
-For PixPro, DeCon-SL is pre-trained using a LARS optimizer for 800 epochs with a batch size of 512 and a learning rate of 1.0 with a weight decay of 1e-5.
-The pre-trainings are conducted with eight NVIDIA A100 GPUs (40GB each).
-To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec.
-3.4.
-For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with the batch size.
-For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC.
-These pre-trainings are performed with one NVIDIA A6000 GPU with 48GB of memory.
+To evaluate the generalization of our approach to different SSL frameworks, we adapt DenseCL and PixPro frameworks with DeCon-SL by pre-training a ResNet-50 encoder with an FCN decoder on COCO 2017 dataset.
+Optimization follows the base frameworks.
+To evaluate the dependence to the pre-training dataset, we also pre-train a similar DeCon-SL adaptation of SlotCon on REFUGE [41] and ISIC [13] datasets.
+We provide all pre-training details in Supplementary Material Section 2.
 3.4 Evaluation protocol
-We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks.
+We evaluate the different pre-trained models on various downstream tasks.
 Unless stated otherwise, we only transfer the pre-trained encoder for the fine-tuning experiments.
-We initialize the network with a pre-trained backbone and fine-tune end-to-end in all cases.
-All fine-tuning results reported in this paper are averaged over three independent runs.
+We initialize the network with a pre-trained encoder and fine-tune end-to-end in all cases.
+All fine-tuning results reported are averaged over three independent runs.
 The pre-trainings, however, are only run once.
-All fine-tuning experiments related to the same dataset are run on a similar platform with the same hardware.
-We perform object detection and instance segmentation using Detectron2 package [49] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
-Object detection and instance segmentation: We perform object detection and instance segmentation on COCO dataset [31]. We fine-tune a Mask R-CNN with FPN backbone following SlotCon [48].
-Semantic segmentation: We evaluate the semantic segmentation performance of a pre-trained ResNet-50 encoder with an FCN decoder on Pascal VOC [18] and Cityscapes [13] following SlotCon [48] and PixCon [39].
-Semantic segmentation on out-of-domain datasets: We assess generalization on REFUGE [38] and ISIC 2017 [10] datasets using the same ResNet-50-FCN architecture, fine-tuned on 5, 25, and 100% of the training data.
-More details on the training setups and hyperparameters are provided in the Supplementary Material Section 2.
+All fine-tuning experiments related to the same downstream task are run on the same hardware.
+We perform semantic segmentation tasks using mmsegmentation [14] version 0.30.0.
+and other remaining tasks using Detectron2 package [56] version 0.6.
+Object detection and instance segmentation: We fine-tune a Mask R-CNN with a pre-trained ResNet-50 or ConvNeXt-Small backbone on COCO 2017 dataset [35].
+Semantic segmentation: We fine-tune a pre-trained ResNet-50 or ConvNeXt-Small with an FCN decoder on Pascal VOC [21], and ResNet-50 with an FCN decoder on ADE20K [63] and Cityscapes [16] following SlotCon [55] and PixCon [42].
+We also fine-tune an ImageNet-1K-pre-trained ConvNeXt-Small backbone with an UPerNet decoder on ADE20K dataset.
+Panoptic segmentation, keypoint detection, and dense pose estimation: We adapt task-specific Detectron2 configurations based on the SlotCon object detection architecture, and fine-tune on the COCO 2017 dataset for panoptic segmentation and keypoint detection, and on the COCO 2014 dataset for dense pose estimation.
+Generalization to out-of-domain datasets: We fine-tune a ResNet-50-FCN architecture on REFUGE [41] and ISIC 2017 [13] datasets for medical semantic segmentation using 5, 25, and 100% of the training data.
+We also fine-tune a ResNet-50-deeplabv3+ architecture for PlantSeg  [54] semantic segmentation and a Faster R-CNN ResNet-50-FPN architecture for object detection on PlantDoc [49] and Detecting Disease  [47] datasets.
+Fine-tuning on agriculture datasets is performed using 10 and 100% of the datasets.
+More details on the training hyperparameters and target datasets are provided in Supplementary Material Section 2.
 4 Results
 4.1 Experiments
-Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML Small and Large ($\alpha=0$ and $dropout=0.5$) adaptations of SlotCon trained with a ResNet-50 backbone.
-The architecture was pre-trained for 800 epochs on COCO and COCO+ and for 200 epochs on ImageNet-1K, and then fine-tuned on various segmentation and object detection tasks.
-Only the pre-trained encoder was transferred for the fine-tuning.
-Results for PixCon pre-trained on ImageNet-1K are not reported, as the corresponding checkpoint or evaluation was not made available in the original publication.
-All fine-tuning results were averaged over three runs.
-means that we reimplemented the framework, * means that the results are taken from the original paper.
-Pret. Dataset Framework Pret. Dec. Object Detection Instance Segmentation Semantic Segmentation
-COCO COCO VOC City
-AP AP50 AP75 AP AP50 AP75 mIoU mIoU
-COCO PixCon [39]* - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62
-SlotCon [48] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
-DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21
-DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21
-DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45
-COCO+ PixCon [39]* - 41.2 - - 37.1 - - 73.9 77.0
-SlotCon [48] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43
-DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65
-DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00
-ImageNet-1K SlotCon [48] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15
-DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51
-ResNet-50 backbone: Table 1 presents the results of fine-tuning a ResNet-50 backbone pre-trained with the two best pre-training frameworks for dense downstream tasks and DeCon.
-A more comprehensive comparison with additional previously published methods is provided in the Supplementary Material 4.
-DeCon-ML-L is consistently establishing a new SATO for all tasks, except when pre-training is performed on COCO dataset for Pascal VOC and Cityscapes semantic segmentation fine-tuning, where it reduces the gap between PixCon and SlotCon.
-This supports the fact that the DeCon adpatation is robust across datasets.
-Secondly, we can see that all of DeCon versions consistently improve upon the base SlotCon framework for any of the pre-training datasets used.
-Notably, DeCon-ML-S, whose parameter count is similar to SlotCon, exhibits better performance than SlotCon.
-Table 2: Performance of DeCon-ML-L ($\alpha=0$ and $dropout=0.5$) adaptation of SlotCon using a ConvNeXt-small encoder.
-The architecture was pre-trained for 800 epochs on COCO, and then fine-tuned on various segmentation and object detection tasks.
-All fine-tuning results were averaged over three runs.
+Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML (Small/Large, $\alpha=0$, dropout=0.5) adaptations of SlotCon with a ResNet-50 backbone pre-trained for 800 epochs on COCO/COCO+ and 200 epochs on ImageNet-1K, then fine-tuned on different downstream tasks.
+Only pre-trained encoders were transferred.
+PixCon ImageNet-1K results are not reported due to unavailable checkpoints.
+Results are averaged over three runs.
+Re-implemented, * Results from original paper, $\lozenge$ Result from SlotCon [55] paper.
+Pret. Dataset Framework Backbone Pret. Dec. Object Det. Instance Seg. Semantic Seg. Dense pose Est. Panoptic Seg. Keypoint Det.
+COCO COCO VOC City ADE COCO COCO COCO
+AP AP50 AP75 AP AP50 AP75 mIoU mIoU mIoU GPS AP GPSm AP PQ AP
+[…]
+COCO PixCon (2024) [42]* ResNet-50 - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62 38.00 - - - -
+SoCo-D (2024) [44]* - 40.3 60.1 44.0 35.1 56.9 37.6 - - - - - - -
+SlotCon (2022) [55] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95 38.57 63.39 64.84 40.31 65.66
+DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21 38.81 63.50 64.98 40.44 65.77
+DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21 38.36 63.75 65.08 40.52 65.66
+DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45 38.70 63.72 64.94 40.90 65.88
+COCO+ PixCon (2024) [42]* ResNet-50 - 41.2 - - 37.1 - - 73.9 77.00 38.80 - - - -
+SlotCon (2022) [55] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43 39.11 - - - -
+DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65 39.25 - - - -
+DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00 39.04 - - - -
+ImageNet-1K Supervised $\lozenge$ ResNet-50 - 39.7 59.5 43.3 35.9 56.6 38.6 74.4 74.6 37.9 - - - -
+DINO (2021) [6] - 40.24 60.25 44.13 36.47 57.49 39.20 73.09 75.57 37.30 - - - -
+SlotCon (2022) [55] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15 38.97 - - - -
+DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51 39.01 - - - -
+Table 2: Performance of DeCon-ML-L ($\alpha=0$, dropout = 0.5) and SlotCon with ConvNeXt-S encoder, pre-trained for 800 epochs on COCO and fine-tuned on segmentation and detection tasks.
 Pret. Dataset Framework Pret. Dec. Obj. Det. Inst. Seg. Sem. Seg.
@@
 DeCon-ML-L FPN 44.71 40.37 73.81
-ConvNeXt backbone: Table 2 presents the results of the fine-tuning of models pre-trained with a ConvNeXt-small backbone using DeCon-ML-L and the original SlotCon.
+Table 3: Comparison of ViT and ConvNeXt models pre-trained on ImageNet-1K.
+We pre-trained a ConvNeXt-S encoder using DeCon-SL (FCN decoder, $\alpha=0.25$).
+All models are fine-tuned with an UPerNet decoder [58] on ADE20K.
+*Results from MixedAE [7]; †Results using two decoder levels to match with DeCon-SL decoder.
+Framework Pret. Backb. Backb. Params. (M) Pret. Epoch ADE20k (mIoU)
+MoCov3 (2021) [12]* ViT-B 88 600 46.8
+DINO (2021) [6]* ViT-B 88 1600 46.9
+BEiT (2021) [2]* ViT-B 88 300 44.7
+MAE (2022) [24]* ViT-B 88 300 46.7
+MixedAE (2023) [7]*† ViT-B 88 300 47.4
+DeCon-SL [ours] ConvNeXt-S 50 250 48.02
+ResNet-50 backbone: Table 1 and Supplementary Material Table S1 present the results of fine-tuning a ResNet-50 backbone pre-trained with various contrastive SSL frameworks.
+We mainly compare with SlotCon and PixCon as they demonstrated the most competitive performance in dense downstream tasks among previously published contrastive methods.
+DeCon-ML-L is consistently establishing a new SOTA for all tasks when pre-trained on ImageNet-1K and is consistently improving upon its base framework SlotCon and showing SOTA results for most of the evaluated downstream tasks when pre-trained on COCO and COCO+ dataset.
+Notably, it reduces the gap between PixCon and SlotCon for Pascal VOC and Cityscapes semantic segmentation when pre-trained on COCO.
+This supports the fact that the DeCon adaptation is robust across pre-training datasets and fine-tuning tasks.
+Moreover, DeCon-ML-S and DeCon-SL, whose training costs are comparable to SlotCon exhibit better performance than SlotCon and other contrastive frameworks.
+Finally, Tab.
+1 also shows that performance improvement compared to SlotCon are robust to other downstream tasks such as dense pose estimation, panoptic segmentation and keypoint detection.
+ConvNeXt backbone and comparisons with ViT: Table 2 shows the fine-tuning performance of a ConvNeXt-Small backbone pre-trained using DeCon-ML-L and the base SlotCon framework.
 A consistent improvement is seen with our encoder-decoder framework, confirming the robustness of our method to more modern and bigger backbones.
-Notably, the performance of the model as well as the performance improvement appear larger than for a smaller ResNet-50 backbone.
-Table 3: Performance of DeCon-SL ($\alpha=0.5$) adaptation of DenseCL and PixPro using a ResNet-50 encoder.
-The architectures were pre-trained for 800 epochs on COCO, and then fine-tuned on Pascal VOC and Cityscapes semantic segmentation.
-All fine-tuning results were averaged over three runs.
+Notably, both the absolute performance and the relative improvement are greater than those observed with the ResNet-50 backbone.
+Table 3 presents a performance comparison between a ConvNeXt-S model pre-trained on ImageNet-1K with DeCon-SL and ViT backbones pre-trained using various frameworks.
+Despite its smaller size and shorter pre-training schedule, the ConvNeXt-S backbone, pre-trained for dense tasks, outperforms all competing methods.
+Table 4: Performance of DeCon-SL ($\alpha=0.5$) adapting DenseCL and PixPro with a ResNet-50 encoder, pre-trained for 800 epochs on COCO and fine-tuned on Pascal VOC and Cityscapes.
 Pret. Dataset Framework Pret. Dec. Transfer Sem. Seg.
@@
 Enc + Dec 73.08 76.19
-Generalizability across frameworks: Table 3 presents semantic segmentation fine-tuning results of models pre-trained with DenseCL and PixPro, and their DeCon-SL adaptation.
-The performance of the model pre-trained with a decoder is improved compared to the base framework and improves even further when the pre-trained decoder is transferred, supporting the hypothesis that DeCon could be adapted to multiple frameworks.
-Transferring the pre-trained decoder however, is not always beneficial to downstream performance in the case of SlotCon.
-More details can be found in Supplementary Material 5.
-Out-of-domain tasks: Table 4 depicts the fine-tuning performance of a model pre-trained on COCO to out-of-domain datasets.
-We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased.
-Further improvement was achieved when transferring the pre-trained decoder along with the encoder for ISIC dataset.
+Generalizability across frameworks: Table 4 presents semantic segmentation fine-tuning results for models pre-trained with DenseCL and PixPro, along with their DeCon-SL adaptations.
+Pre-training with a decoder improves performance over the base frameworks, and transferring the decoder further enhances results, supporting the hypothesis that DeCon can be adapted to multiple frameworks.
+However, transferring the pre-trained decoder does not consistently benefit downstream performance in the case of SlotCon, as seen in Tab.
+5 and Supplementary Material 5.
+Out-of-domain tasks: Table 5 shows fine-tuning results of COCO-pre-trained encoders on REFUGE and ISIC.
+DeCon-SL pre-training consistently outperforms SlotCon, with larger gains under limited data.
+Transferring both encoder and decoder further improves performance on ISIC.
 Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset.
 However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training.
-In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if only the encoder was transferred.
-We also observed that pre-trained models outperformed models trained with randomly initialized weights.
-This gap was more notable in REFUGE dataset (over 20%).
-Table 4: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights.
-We pre-trained SlotCon and its DeCon-SL adaptation with a ResNet-50 encoder and an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets.
-We fine-tuned on REFUGE and ISIC datasets under various dataset sizes.
-We averaged the fine-tuning mIoU results over three runs.
+In both ISIC and REFUGE pre-trainings, DeCon-SL outperformed SlotCon.
+Table 6 shows that DeCon variants also perform better than SlotCon in agricultural object detection and semantic segmentation tasks using 10 or 100% of the datasets.
+DeCon pre-trained models also outperformed randomly initialized ones.
+Table 5: Transfer performance with SlotCon and DeCon-SL (ResNet-50, FCN, $\alpha{=}0.5$) pre-trained on COCO, REFUGE, and ISIC, then fine-tuned on out-of-domain medical datasets (REFUGE and ISIC sem.
+seg.
+) in various data settings.
 SSL strategy Loss Transfer Pre-trained on COCO Pre-trained on Downstream Data
 REFUGE ISIC REFUGE ISIC
+mIoU mIoU mIoU mIoU mIoU mIoU mIoU mIoU
 $L_{enc}$ $L_{dec}$ enc dec 5% 25% 100% 5% 25% 100% 100% 100%
@@
 72.09 77.27 83.25 76.00 79.97 82.94 81.01 83.25
+Table 6: ResNet-50 backbone transfer performance of SlotCon, DeCon-SL (FCN decoder) and DeCon-ML (FPN decoder), pre-trained on COCO and fine-tuned on agriculture datasets with 10% and 100% labeled data.
+Framework PlantDoc (Obj. Det.) Detecting Diseases (Obj. Det.) PlantSeg (Sem. Seg.)
+AP AP AP AP mIoU mIoU
+10% 100% 10% 100% 10% 100%
+Random Init. 6.59 19.81 19.14 34.76 16.96 24.96
+SlotCon 17.53 38.37 26.54 48.45 20.72 28.78
+DeCon-SL ($\alpha=0.25$) 17.06 39.94 26.81 48.85 20.56 29.39
+DeCon-ML-S ($\alpha=0$, drop=0.5) 17.82 39.40 26.82 48.53 21.00 29.16
+DeCon-ML-L ($\alpha=0$, drop=0.5) 17.98 40.03 27.16 48.57 20.66 29.77
 4.2 Ablation studies
 Table 7 presents an ablation study of the components in our DeCon-ML-L framework.
-The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, even when only the encoder is transferred, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision.
+The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision.
 This combination yields the most substantial gains, highlighting the critical role of channel dropout when paired with deep supervision in our approach.
 Table 7: Ablation of the different components of DeCon-ML-L and their impact on the downstream performance.
-We averaged all fine-tuning results over three runs.
 $\Delta$ is the performance gap between the previous framework and the framework updated with the corresponding new component.
@@
 +Enc-Dec Loss Weight ($\alpha$) Tuning (DeCon-ML-L) 41.18 +0.08 37.12 +0.07
-Table 6 shows DeCon-ML performance with different values of encoder loss weight $\alpha$ and channel dropout.
+Table 8 shows DeCon-ML performance with different values of encoder loss weight $\alpha$, channel dropout, and number of decoder levels.
 A channel dropout probability of 0.5 provided the best downstream performance for DeCon-ML-L.
@@
 An $\alpha$ value of 0 gave the best downstream performance meaning that the decoder loss can replace the encoder one and pre-train efficiently an encoder on its own.
+We also ablated with number of decoder levels in DeCon-ML-L where 4 decoder levels yielded the best averaged performance across tasks.
 For DeCon-SL, an $\alpha$ value of 0.25 resulted in the best downstream performance, suggesting that having an encoder loss at the bottleneck is still beneficial to the pre-training in the absence of skip connections.
-Table 6: Impact of the channel dropout probability and of the encoder loss weight $\alpha$ on DeCon.
-We pre-trained DeCon adaptations of SlotCon framework on COCO dataset using different parameters and evaluate the pre-trained encoder fine-tuning on downstream tasks that use the same decoder architecture.
-We averaged all fine-tuning results over three runs.
-Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Frame work Deco der $\alpha$ Sem. Seg.
+Table 8: Ablation of channel dropout, encoder loss weight ($\alpha$), and decoder levels in DeCon, pre-trained on COCO for 800 epochs and evaluated by fine-tuning on downstream tasks.
+Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Abla tion Frame work Deco der $\alpha$ Dec Levels Sem. Seg.
 COCO COCO VOC City
 AP AP mIoU mIoU
-Dropout
-DeCon-ML-L
-FPN
-0.50 0 40.85 36.83 - - - - -
-0.25 40.95 36.97 - - - - -
-0.50 41.10 37.05 - - - - -
-0.75 41.02 37.05 - - - - -
-$\alpha$
-0.75 0.50 40.98 36.90
-DeCon-SL (No Drop)
-FCN
-0.75 71.98 76.18
-0.50 41.10 37.05 0.50 72.42 75.62
-0.25 41.05 37.13 0.25 73.01 76.21
-0 41.18 37.11 0 72.17 75.71
+Dropout DeCon-ML-L (4 decoder levels) FPN 0.50 0 40.85 36.83 Dec. Level DeCon-ML-L (Drop = 0.5) FPN 0 2 72.70 76.07
+0.25 40.95 36.97 3 73.14 76.10
+0.50 41.10 37.05 4 72.92 76.45
+0.75 41.02 37.05 - - - -
+$\alpha$ 0.75 0.50 40.98 36.90 $\alpha$ DeCon-SL (No Drop) FCN 0.75 - 71.98 76.18
+0.50 41.10 37.05 0.50 - 72.42 75.62
+0.25 41.05 37.13 0.25 - 73.01 76.21
+0 41.18 37.11 0 - 72.17 75.71
 5 Discussion
-Table 7: COCO and COCO+ pretraining downstream performance improvement with DeCon-ML-L and Cohen’s d value. PixCon performance was extracted from the original paper.
-Method COCO obj det. COCO Inst. Seg. VOC Sem. Seg. City Sem. Seg.
-$\Delta$AP $\Delta$AP $\Delta$mIoU $\Delta$mIoU
-COCO Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
-SlotCon (2022) (Comp. to prev. works) +0.50 +0.40 -0.40 +0.10
-PixCon (2024) (Comp. to SlotCon) 0.00 +0.06 +1.30 +0.51
-DeCon-ML-L (Comp. to SlotCon) +0.37 +0.32 +1.42 +0.50
-Cohen’s d (DeCon-ML-L VS SlotCon) 2.35 1.94 6.37 2.51
-COCO+ Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
-SlotCon (2022) (Comp. to prev. works) +0.60 +0.50 Not avail. Not avail.
-PixCon (2024) (Comp. to SlotCon) -0.50 -0.50 -0.20 +0.40
-DeCon-ML-L (Comp. to SlotCon) +0.45 +0.27 +1.43 +0.57
-Cohen’s d (DeCon-ML-L VS SlotCon) 5.42 3.21 5.25 1.22
 In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework.
-We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation when pre-trained on COCO dataset, and consistently outperforms previous methods when pre-trained on COCO+, and ImageNet except for Cityscapes semantic segmentation (Tab.
+We showed that our proposed SSL adaptation achieved SOTA performance in object detection, instance and semantic segmentation tasks when pre-trained on COCO+ and ImageNet-1K datasets, and in most tasks when pre-trained on COCO (Tab.
 1).
-This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab.
-1).
+This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, even in tasks beyond object detection, instance and semantic segmentation, regardless of the homogeneity between pre-training and fine-tuning decoders.
+Figure 2: DeCon-ML-L and SlotCon pre-training loss dynamics.
 We introduced a combined loss where encoder and decoder losses contribute to the total loss in a weighted manner.
-Figure 2: Pre-training loss dynamics for DeCon-ML-L and SlotCon.
-Figure 2 illustrates the loss behavior across different pre-trainings when both the encoder and decoder are used.
-As shown, decoder pre-training does not influence the encoder’s loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing.
-We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks.
-This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28].
+Figure 2 illustrates losses behavior across different pre-trainings when both the encoder and decoder are used.
+As shown, decoder pre-training does not affect the encoder loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing.
+We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models.
+This could be attributed to the non-competing nature of the encoder and decoder losses compared to approaches that use a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [31].
 Additionally, Fig.
@@
 It demonstrates that the decoder learns features that capture similar concepts as the encoder’s, while being more spatially precise.
-Figure 3: Slots, as defined in SlotCon, learned at different outputs of our architecture.
+Figure 3: Slots, as defined in SlotCon [55], learned at different outputs of our architecture.
 The top left image from COCO validation dataset is overlayed with the slot that was the most represented in original SlotCon’s encoder output feature map, resized to the image’s shape.
@@
 Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML).
-Framework specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (shown in Supplementary Material Section 8).
-We demonstrated that our method is effective with a different type of encoders.
+Framework-specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (Supplementary Material Section 8).
+We demonstrated that our method is effective with different encoders.
 Specifically, pre-training a ConvNeXt-Small backbone, which is both more efficient and larger than the ResNet-50 backbone, also resulted in superior downstream performance when integrated with a decoder in the DeCon-ML-L framework.
 These results attest to the scalability and practical utility of the proposed approach for pre-training modern backbone architectures.
+Additionally, Tab.
+3 shows that a ConvNeXt-S backbone pre-trained with DeCon-SL for dense tasks outperforms larger ViT-based SSL methods, despite using fewer pre-training epochs.
 DeCon was also used to successfully adapt three different SSL frameworks using different concepts (pixel level VS object/group level) to learn local information (Tab.
-3).
-This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods designed for dense downstream tasks.
-Table 8: Pre-training parameter count, GPU setup, and time required per epoch for the different architectures.
+4).
+This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods targeting dense tasks.
+Table 9: Pre-training parameter count and computational cost for the different architectures.
 We study various pre-training setups on the COCO dataset using the SlotCon framework.
 Module sizes are given in millions of parameters.
-Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU setup Time per epoch (s)
-SlotCon 23.51 x 2 0 71.51 0 118.52 1 * A100 80GB 255
-DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1 * A100 80GB 300
-DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1 * A100 80GB 280
-DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 8 * H100 80GB 165
-DeCon adaptation of an SSL framework provided notable advantages for fine-tuning on small-scale datasets and out-of-domain dense prediction tasks, outperforming the encoder-only baseline (see Tab.
-4).
-This improvement persisted even when pre-training was conducted directly on the small target datasets.
-Furthermore, DeCon achieved superior performance compared to related frameworks when only a limited fraction (5–25%) of the original labeled data was available for fine-tuning.
-These results highlight the value of DeCon in settings with very limited annotated samples or small-scale, domain-specific datasets, such as those in medical or agricultural applications, regardless of whether the pre-trained decoder is transferred.
+GPU-days are calculated on H100 80GB GPUs.
+Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU- days
+SlotCon 23.51 x 2 0 71.51 0 118.52 1.53
+DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1.77
+DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1.78
+DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 11.00
+DeCon adaptation of SSL frameworks yields clear advantages on small-scale datasets and out-of-domain dense prediction tasks, outperforming encoder-only baselines (Tab.
+5, Tab.
+6).
+These gains persist even when pre-training on the small target datasets, or when fine-tuning with only a fraction of labeled data.
+This highlights DeCon’s value in domains with limited annotations, like medical and agricultural applications.
 Performance of deep learning models can vary across multiple runs due to inherent stochasticity.
 To mitigate the impact of randomness while considering computational resource constraints, each experiment was conducted three times, with all fine-tunings for a specific downstream task performed on similar hardware.
-The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8> 0.8) in Tab.
-7, exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests.
-Details on the computation of these values are provided in Supplementary Material 10.
-Furthermore, the performance gains shown in Tab.
-7 are comparable to those reported in prior studies.
-One limitation of our approach is its high memory and computational demand.
-However, we have shown in Tab.
-1 that if GPU is a constraint, the user can decide to reduce the number of decoder levels used in the pre-training and tune the decoder projector layers to reduce the parameter count of DeCon-ML architecture and match the parameter count of the original framework (as shown in Tab.
-8), while still improving the performance of the pre-training.
+The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8) (in Supplementary Material Tab.
+S6), exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests.
+Details on the computation of these values are provided in Supplementary Material Section 10.
+Furthermore, the performance gains, shown in Supplementary Material Tab.
+S6, are comparable to those reported in prior studies.
+A limitation of our approach is its high memory and computational demand.
+However, as shown in Tab.
+1, when GPU access is limited, reducing the number of decoder levels and tuning the decoder projector layers can decrease the parameter count and training cost of DeCon-ML (e.g., DeCon-ML-S in Tab.
+9) while still enhancing pre-training performance.
+It is also notable that DeCon-SL and DeCon-ML-S have comparable training costs as their baseline.
 Brempong et al.
@@
 Future works could involve extending our proposed DeCon adaptation to offer a multi-stage continual pre-training in a contrastive manner.
+We also plan to adapt DeCon to ViT architectures in the future.
 6 Conclusion
-In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation.
+In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of SSL approaches for various dense predictions tasks.
 We showed that this adaptation improves the encoder’s representation significantly.
@@
 Acknowledgments
+We extend our sincere gratitude to the University of Calgary VPR Office for their generous support through VPR Catalyst Grants, which was instrumental in making this research possible.
+We also acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC Discovery Grant RGPIN-2024-04966) and Alberta Innovates.
+This work was partly supported by Alberta Innovates Graduate Student Scholarship and T.
+Chen Fong Doctoral Research Excellence Scholarship in Imaging Science.
 Computing resources were provided by the University of British Columbia Cloud Innovation Centre, powered by Amazon Web Services https://cic.ubc.ca/.